### v1

# Practical Collaborative Perception: A Framework for Asynchronous and Multi-Agent 3D Object Detection

Abstract
In this paper, we improve the single-vehicle 3D object detection models using LiDAR by extending their capacity to process point cloud sequences instead of individual point clouds. In this step, we extend our previous work [1] on rectification of the shadow effect in the concatenation of point clouds to boost the detection accuracy of multi-frame detection models. Our extension includes incorporating HD Map and distilling an Oracle model. Next, we further increase the performance of single-vehicle perception using multi-agent collaboration via Vehicle-to-everything (V2X) communication. We devise a simple yet effective collaboration method that achieves better bandwidth-performance tradeoffs than prior arts while minimizing changes made to single-vehicle detection models and assumptions on inter-agent synchronization. Experiments on the V2X-Sim dataset show that our collaboration method achieves 98% performance of the early collaboration while consuming the equivalent amount of bandwidth usage of late collaboration which is 0.03% of early collaboration. The code will be released at https://github.com/quan-dao/practical-collab-perception
Index Terms:
V2X Collaborative Perception, 3D Object Detection, Scene Flow.
I Introduction
Object detection, which amounts to localizing objects as 3D bounding boxes, is a fundamental autonomous driving software stack module. It provides input to downstream tasks such as object tracking, motion forecast, and navigation. Due to the need for accurately localizing objects in 3D, state-of-the-art methods use LiDARs as the primary sensing modality. While significant advancements have been made, two challenges continuously plague LiDAR-based detection models: occlusion and sparsity, especially at long ranges.
As occluded objects in one frame can be visible in others due to their motion and/or the ego vehicle’s motion, and sparse regions can become dense as the ego vehicle approaches, the utilization of point cloud sequences is an appealing candidate for boosting the boundary of 3D object detection. First introduced in [2] as a means for alleviating the extreme sparsity of point clouds obtained by a 32-beam LiDAR, the concatenation of point cloud sequences using ego vehicle pose; thus the name Ego Motion Compensation (EMC), is by far the most popular multi-frame method due to its simplicity. Yet, being highly effective [3, 4, 5, 6, 7]. More recent works such as 3D-MAN [8], MoDAR [9], and MPPNet [10] are built upon this concept by dividing the input sequence into several subsequences whose point clouds are concatenated to produce preliminary detections. These preliminary detections are then fused to result in the final detections.
The major drawback of EMC-concatenated point clouds is that their formulation only accounts for the motion of the ego vehicle while ignoring the motion of dynamic objects, thus their instantiating at multiple locations, which is referred to use the shadow effect [11], during the observation horizon as can be seen in Fig.1. This effect, which is essentially a misalignment in 3D space between objects’ points and their location, results in a misalignment in feature spaces such as on the bird-eye view image, limiting the performance gain brought by an increased number of points (thanks to point cloud concatenation) to stationary and slow-moving objects [8]. In our previous work [1], we resolved this alignment explicitly by rectifying concatenated point clouds using estimated point-wise scene flow. The add-on module named Aligner, which we proposed, achieved competitive scene flow performance while substantially improving the detection accuracy, especially of large objects with a severe shadow effect.
Figure 1: A dynamic car’s appearance on the concatenation of a 10-point-cloud sequence spanning 0.5 seconds. Points are colour-coded according to their time lag with respect to the present.
Despite the power of point cloud sequences, the issue of occlusion and sparsity at long range persists as several regions remain invisible to the ego vehicle due to the geometry of the scene (e.g., a large object obstructs the ego vehicle’s field of view) or the traffic rule that limits the ego vehicle’s motion (e.g., the ego vehicle must stop at a red light). These issues can be safety critical when the ego vehicle navigates complex traffic, such as at intersections where a Waymo report indicates that a quarter of severe accidents involving their autonomous vehicles are due to occlusion [12]. Therefore, in this work, we aim to boost further object detection made by a single-vehicle operating on point cloud sequences using the Vehicle-to-everything (V2X) collaborative perception.
In the collaborative perception framework via V2X, Intelligent Road-Side Units (IRSU), which are equipped with advanced sensing systems (e.g., comparable to an autonomous vehicle) and located at elevated locations that maximize their field of view, and other vehicles can make measurements in regions that are inaccessible or partially occluded to others. Together they can come up with a complete representation of the environment. The enhanced perception capacity, thanks to V2X, comes with several new technical challenges; the most notorious among them is the performance-bandwidth tradeoff which revolves around two questions (i) which information to be broadcasted and (ii) how to fuse the exchanged information.
This tradeoff establishes a design spectrum ranging from early collaboration, which exchanges raw sensory measurements - point clouds in the context of this paper, thus consuming the maximum bandwidth, to achieve the maximal performance to late collaboration that exchanges outputs (e.g., object detection as 3D bounding boxes) to achieves minimal bandwidth usage at the expense of limited performance. Naturally, researches on V2X collaboration frameworks [13, 14, 15, 16] are drawn toward the middle of this spectrum, thus the category’s name of mid-collaboration, where intermediate representations such as bird-eye view (BEV) images of agents’ surrounding environment are chosen for broadcasting in an attempt of balancing the two mutually excluding design targets. While the motivation is just, most mid-collaboration methods require making substantial changes to the architecture of single-agent perception models to accommodate the fusion module where the combination of exchanged representations takes place. More importantly, most methods make strong assumptions about the synchronization among connected agents. For example, DiscoNet [14] considers a perfectly synced setting where agents share the same clock, collect and process point clouds at the same rate and the transmission/ receiving of BEV images experience zero latency. V2VNet [13] and ViT-V2X [15] account for latency by postulating a global misalignment between exchanged BEV images and the BEV image of the ego vehicle caused by the movement of the ego vehicle between the time step it makes the requests for BEV images and the time step the exchanged BEV images are received. This means these works implicitly assume agents in the V2X network obtain point clouds synchronously.
Taking a different approach to mid-collaboration, we aim for a more practical collaboration framework that emphasizes minimal
  • Bandwidth consumption
  • Changes made to single-agent models
  • Inter-agent synchronization assumption
We ground our design at minimal bandwidth consumption which can only be achieved by exchanging agents’ output which detected objects. This design choice naturally satisfies the second design target as it dismisses the need for complex mid-representation fusion modules. We achieve the final target by only assuming that connected agents share a common time reference which is practically achievable using GPS time.
The challenge posed by our relaxed inter-agent synchronization assumption is that information (i.e., detected objects) broadcasted by agents in the V2X network may never have the same timestamp as the query made by the ego vehicle. In other words, the detections made by other agents that are available on the V2X network always have an older timestamp compared to the current timestamp of the ego vehicle. This timestamp mismatch results in a misalignment between exchanged detected objects and their associated ground truths (if the detections are true positive). We resolve this misalignment by predicting objects’ velocities simultaneously with their poses using the Aligner module developed by our previous work [1]. Specifically, each detected object pools point-wise scene flow produced by the Aligner to compute its velocity, which is then used to propagate this object from one timestamp to another. As the scene flow plays a more critical role in the V2X collaborative perception, we first extend our Aligner by incorporating HD Maps into its framework and improving its supervision by distilling an Oracle model (details are in Sec.III-A2). The resulting module is dubbed Aligner++.
Aware of the limitation of the late collaboration whose details are presented in Sec.II, we derive a novel collaboration strategy that fuses detections made by connected agents at a past timestamp with the current point cloud acquired by the ego vehicle using inspiration from the recent advance in multi-frame single-vehicle object detection [9].
This paper makes the following contributions:
  • Extending our previous work [1] to further boosts single-vehicle object detection accuracy and achieves more accurate scene flow prediction
  • Deriving a practical framework for V2X collaborative perception that consumes minimal bandwidth while achieving competitive performance compared to the early collaboration in the perfect synchronization setting
  • Performing extensive evaluations on NuScenes [2], KITTI [17], and V2X-Sim [18] dataset to verify our method
II Related Works
A V2X cooperative perception framework comprises two main components connected autonomous vehicles (CAVs) and intelligent roadside units (IRSUs), both equipped with advanced sensing systems to perceive their environment. IRSUs are stationery installed at elevated positions to obtain minimally occluded fields of view. CAVs and IRSUs can exchange information to enhance their perception capacity (e.g., seeing beyond occlusion). In the context of this work, the sensing system of CAVs and IRSUs is restricted to a single LiDAR and the perception task focuses on 3D object detection.
As described in the previous section, the main challenge of V2X cooperative perception is the performance-bandwidth tradeoff which establishes a design spectrum ranging from early to late collaboration. In the early collaboration approach, as depicted in Fig.2, agents exchange their raw measurements - point clouds. At every timestep, the ego vehicle concatenates its own point cloud with those obtained by other CAVs and IRSUs to form the input for its perception module. This combined point cloud offers a comprehensive view of the scene with minimal occlusion and sparsity thanks to the diverse perspectives of connected agents, thus being regarded as the upper bound of the performance of the V2X collaborative perception [18, 19]. However, due to the significant amount of bandwidth required to transmit raw point clouds (the order of 10 MB), the early collaboration strategy is not feasible for real-world deployments.
Figure 2: Architecture of the early collaboration.
On the other extreme of the performance-bandwidth tradeoff, the late collaboration approach focuses on minimizing bandwidth usage by exchanging only the agents’ predictions, specifically the detection results in the form of 3D bounding boxes. As illustrated in Fig.3, each agent independently conducts object detection on its own point cloud. Subsequently, the agent merges its own predictions with those made by others to generate the final output. While this strategy is more feasible for real-world deployments thanks to its minimal bandwidth consumption, it exhibits significantly lower performance gains compared to early fusion. The limited interaction among agents in the late collaboration approach contributes to this inferior performance. In noisy environments where latency is a factor, the late collaboration strategy even underperforms single-vehicle perception [15].
Figure 3: Architecture of the late collaboration.
The mid-collaboration strategy aims to find a balance between performance and bandwidth consumption by exchanging intermediate scene representations generated by the backbone of the agents’ perception model. The motivation behind this approach is that the intermediate scene representation contains more contextual information compared to the final output (i.e., 3D bounding boxes), enabling deeper interaction among agents. Moreover, this representation is more compact than raw point clouds since it has been reduced in size through a series of convolution layers in the backbone. It can also be further compressed using an autoencoder to minimize bandwidth usage. While the idea is elegant, implementing the intermediate collaboration strategy requires a range of modules, shown in Figure 4, including compressor, decompressor, and representation fusion, among others, to match the performance of early collaboration. The fusion, in particular, is quite intricate as it involves learnable collaboration graphs using techniques such as Graph Neural Networks [13, 14] or Transformers [15] to effectively fuse exchanged representations. Furthermore, dedicated modules are required to account for different practical challenges. For example, [13, 15] use the Spatial Transformer [20] to resolve the global misalignment between the ego vehicle’s representation and others’ caused by the ego vehicle’s motion between when it makes the query and when it receives exchanged information. In [16], the Fused Axial Attention [21] is used to bridge the domain gap between representations made by different detection models (e.g., PointPillar [22] and VoxelNet[23]) used by different agents in the V2X network. More importantly, most mid-collaboration methods make strong assumptions about inter-agent synchronization which is either (i) perfect synchronization where exchanged representations always share the same timestamp [14] or (ii) synchronized point cloud acquisitions, meaning all agents obtain and process point clouds at the same rate and at the same time [13, 15]. Because of these complexities, the real-world deployment of intermediate fusion remains challenging.
Figure 4: V2X Cooperative Perception via the Mid-fusion architecture.
In this work, we aim to resolve the aforementioned complexities of mid-collaboration to obtain a practical framework for V2X collaborative perception. Our design is grounded in minimal bandwidth consumption as this is nonnegotiable in real-world deployment. To achieve this design goal, we choose object detection, in the form of 3D bounding boxes, as the information to be exchanged, which is similar to the late collaboration strategy. We further relax the assumption on inter-agent synchronization to agents sharing a common time reference (e.g., GPS time) and acknowledge that agents produce detections at different rates. As a result, exchanged detections always have older timestamps compared to the timestamp of the query made by the ego vehicle, thus risking a spatial misalignment between exchanged detections and their associated ground truth (if detections are true positive). We resolve this misalignment by predicting objects’ velocity simultaneously with their locations by pooling point-wise scene flow which can be produced by integrating our Aligner++ module to any off-the-shelf BEV-based object detectors. Finally, we avoid the inferior performance of late collaboration by devising a new collaboration strategy that fuses exchanged detections with the ego vehicle’s raw point cloud for subsequent processing using its detection model rather than fusing exchanged detections with detections made by the ego vehicle. An illustration of our approach is presented in Fig.5.
The innovation of our collaboration strategy lies in recognizing the similarity between object detection using point cloud sequences and collaborative detection. In both cases, there is a need to fuse information obtained from diverse perspectives. Point cloud sequences involve capturing the motion of the ego vehicle, which results in varying viewpoints, while collaborative detection entails incorporating insights from other agents present in the environment. By drawing this parallel, we leverage the shared principle of fusing information from multiple perspectives to enhance the accuracy and robustness of both object detection approaches
Figure 5: Our V2X Cooperative Perception.
III Methodology
This section first presents the extension we introduce to our previous work [1] to improve the scene flow estimation as well as further boost the performance gain when integrating our module to single-vehicle object detection models. Next, our strategy for V2X collaborative perception that enables using asynchronous exchanged information is derived.
III-A Improving Single-Vehicle Perception
As mentioned in Sec.I, the accuracy of single-vehicle detection can be significantly improved by using the concatenation point cloud sequences as inputs. However, such concatenation is done by the Ego Motion Compensation method which only accounts for the motion of the ego vehicle resulting in the shadow effect (Fig.1) in the concatenated point cloud. This effect, which is a misalignment in the 3D space between objects’ points and their locations, leads to a misalignment in feature spaces (e.g., BEV representations), thus degrading detection accuracy for dynamic objects. In this section, we recall the approach we use in [1] to handle this effect and present the extensions we make to it.
III-A1 Shadow Effect: Cause and Rectification
Let $\mathcal{P}^{t}=\left\{\mathbf{p}_{j}^{t}=[x,y,z,r,\Delta]~{}|~{}j=1,\dots,N\right\}$ denote a point cloud collected by the ego vehicle at time step $t$. Here, each point $\mathbf{p}_{j}$ has two features namely reflectant $r$ and time-lag $\Delta$ with respect to a predefined time step. The point cloud is expressed in the ego vehicle frame $\mathcal{E}(t)$ measured with respect to a global frame $\mathcal{G}$. To concatenate or a point cloud sequence $\mathcal{S}=\left\{\mathcal{P}^{t-K},\mathcal{P}^{t-K+1},\dots,\mathcal{P}^{t}\right\}$ of length $K+1$, each point of point cloud $\mathbf{P}^{t-\Delta t}$ is transformed to the global frame using the ego vehicle pose at the same time step as in (1), hence the method’s name Ego Motion Compensation (EMC).
$^{\mathcal{G}}\mathbf{p}_{i}^{t-\Delta t}=~{}^{\mathcal{G}}\mathbf{T}_{\mathcal{E}(t-\Delta t)}\cdot~{}\mathbf{p}_{i}^{t-\Delta t}$ (1)
Here, ${}^{d}\mathbf{T}_{s}\in\mathbf{SE(3)}$ represents the rigid transformation that maps points in frame $s$ to frame $d$. EMC undoes the motion of the ego vehicle; however, it fails to account for objects’ motion. As a result, the appearance of a dynamic object in the concatenated point cloud comprises several instances, each corresponding to the object’s poses in the global frame at a particular time step, as in Fig.1.
Such distortion can be rectified by first transforming an object’s points collected at different time steps to its body frame, and then from to the global frame at the desired time step. This two-step transformation is illustrated in (2).
$^{\mathcal{G}}\mathbf{\hat{p}}^{t-\Delta t}=~{}^{\mathcal{G}}\mathbf{T}_{\mathcal{O}(t)}\cdot~{}^{\mathcal{O}(t-\Delta t)}\mathbf{T}_{\mathcal{G}}\cdot~{}^{\mathcal{G}}\mathbf{p}^{t-\Delta t}$ (2)
Here, $\mathcal{O}(t)$ represents the object frame at time $t$. The point’s subscript is omitted for clarity.
Equation (2) implies that ${}^{\mathcal{O}(t)}\mathbf{T}_{\mathcal{O}(t-\Delta t)}$ is the identity transformation, which is the case for rigid objects. More importantly, the computation of the rectification transformation, ${}^{\mathcal{G}}\mathbf{T}_{\mathcal{O}(t)}\cdot~{}^{\mathcal{O}(t-\Delta t)}\mathbf{T}_{\mathcal{G}}$, requires object poses which are not available at test time. Our previous work [1] develops a model that estimates such transformation directly from the concatenation of point cloud sequences. The following section briefly recalls the method used in [1] and presents our extension.
III-A2 Learning to Rectify the Shadow Effect
In [1], we propose a drop-in module, named Aligner, for 3D object detection models that create a Bird-Eye View (BEV) representation, $\mathbf{B}\in\mathbb{R}^{C\times H\times W}$, of the scene as an intermediate output. Our module first computes the feature of $\mathbf{f}_{i}\in\mathbb{R}^{C}$ of a point $\mathbf{p}_{i}$ by bilinear interpolating on $\mathbf{B}$ using its projection to the BEV. Each point feature is then decoded into the point’s scene flow $\mathbf{d}_{i}\in\mathbb{R}^{3}$ representing how much this point needs to offset to rectify the shadow effect. After being rectified, the new point cloud is used to scatter the set of point features $\left\{\mathbf{f}_{i}\right\}$ to the BEV to obtain a new BEV image $\mathbf{B}^{\prime}$. While undergoing minimized impact of the shadow effect thanks to the rectification, $\mathbf{B}^{\prime}$ is sparse as only occupied pillars have features, which has a negative effect on detection accuracy [24]. On the other hand, $\mathbf{B}$ is semi-dense thanks to the occupancy leakage caused by the convolutions in the detection models’ backbone but possesses feature misalignment due to the shadow effect. To utilize the best of the two representations, they are fused before being fed to the Region Proposal Networks (RPN) for 3D bounding boxes. The complete pipeline is summarized in Fig.6.
Figure 6: Aligner
The learning target $\mathbf{d}^{*}$ for the scene flow $\mathbf{d}$ of a point $\mathbf{p}^{t-\Delta t}$, which has the timestamp $(t-\Delta t)$ (i.e., collected at time step $t-\Delta t$), is the difference between its location and its rectified location computed by (2).
$\mathbf{d}^{*}=~{}^{\mathcal{G}}\mathbf{\hat{p}}^{t-\Delta t}-~{}^{\mathcal{G}}\mathbf{p}^{t-\Delta t}$ (3)
A 3D bounding box is parameterized by a seven-vector comprised of its center coordinate $\left[c_{x},c_{y},c_{z}\right]$, its size $\left[l,w,h\right]$, and its heading $\theta$. The learning target is defined according to the framework which the RPN follows (e.g., anchor-based [25] or center-based [4]).
To improve the accuracy of [1] in estimating scene flow, thus ultimately improving object detection, we introduce two extensions (i) incorporating HD Map and (ii) distilling an Oracle model. Previous works in integrating HD Maps into 3D object detection models ([26, 27]) opt for a mid-fusion approach that concatenates rasterized maps with backbone-made BEV representations. However, this approach is incompatible with copy-paste data augmentation [25], which randomly samples ground truth objects from a database and pastes them to each point cloud used for training, as pasted objects do not necessarily adhere to the semantics and geometry of the map (e.g., cars appear inside a building). Moreover, they only utilize the map’s binary channels such as drivable areas, sidewalks, or car parks while omitting the lane direction which is potentially helpful for estimating scene flow and objects’ heading direction.
Aware of these two limitations, we propose to extract the map feature for each 3D point and use it to augment the point’s raw features prior to further processing (e.g., building a ground truth database or computing a BEV representation for object detection). The map feature extraction process is illustrated in Fig.7 where points’ coordinate in the map’s BEV is used for nearest neighbor interpolation. This attachment of the map features to points results in an early fusion between HD Maps and LiDARs, rendering the concatenation of maps’ channels to backbone-made BEV representations unnecessary. As a result, HD Maps are made compatible with copy-paste data augmentation.
Figure 7: Extracting map features for points.
Figure 8: Alinger++ ’s Teacher-student framework
To improve the quality of the fused BEV representation, $\mathbf{B}^{f}$ in Fig.6, in terms of minimizing the feature misalignment caused by the shadow effect and reducing the sparsity, we use the teacher-student framework shown in Fig.8. In detail, we use ground truth objects’ trajectories to rectify the concatenation of point cloud sequences and use the shadow-effect-free results, illustrated in Fig.9, to train a teacher model. After the teacher converged, its BEV representation $\mathbf{B}^{t}$ is used to guide the student’s fused BEV representation $\mathbf{B}^{f}$ by optimizing the students’ weight so that the difference measured by the $\mathit{L}_{2}$ loss between these two representations is minimized.
Figure 9: Comparison between concatenation of a point cloud sequence (a) and its rectification using objects’ ground truth trajectories (b). Points are color-coded according to their time-lag with respect to the present. The color bar is shown in (c).
III-B V2X Collaborative Perception Framework
The V2X setting that we target in this paper comprises multiple CAVs and IRSUs which are collectively referred to as agents. An agent $\mathcal{A}_{i}$ is equipped with a LiDAR to localize in a common global frame $\mathcal{G}$ and detect objects in its surrounding environment. The detection is based on the processing of point cloud sequences using a detection model made of the integration of the Aligner++ presented in the Sec.III-A2 into an off-the-shelf single-frame object detector such as PointPillar [22]. At a time step $t_{i}$, agent $\mathcal{A}_{i}$ uses a $K$-point-cloud sequence $\mathcal{S}_{i}=\left\{\mathcal{P}^{t_{i}-K+1}_{i},\dots,\mathcal{P}^{t_{i}}_{i}\right\}$ as an input to its detection model. An object $b_{i,j}$ detected by agent $\mathcal{A}_{i}$ is parameterized by a nine-vector $[x,y,z,w,l,h,\theta,s,c]$. The first seven numbers localize the object by its center location $[x,y,z]$, size $[w,l,h]$, heading direction $\theta$. The last two numbers $s$ and $c$ respectively denote confidence score $s$ and the predicted class $c$.
Upon receiving a query having timestamp $t$, agent $\mathcal{A}_{i}$ will communicate its detection $\mathcal{B}_{i}^{t_{i}}=\left\{\mathbf{b}_{i,j}\right\}_{j=1}^{M_{i}}$ and metadata produced timestamp $t_{i}$ that is prior to and closest to $t$. The agent’s metadata produced at a timestamp $t_{i}$ is made of the timestamp itself and the agent’s pose $\mathcal{E}_{i}(t_{i})$ at this timestamp.
Given this setting, we aim to enhance the ego vehicle’s capacity of detecting objects by fusing its point cloud sequence $\mathcal{S}_{e}=\left\{\mathcal{P}^{t-K+1}_{e},\dots,\mathcal{P}^{t}_{e}\right\}$ with the MoDAR interpretation of predictions $\mathcal{B}_{i}^{t_{i}}$ made by other agents. The following section provides an overview of the origin of MoDAR and presents in detail how we adapt this concept to collaborative perception via the V2X context.
III-B1 MoDAR for Object Detection on Point Cloud Sequences
MoDAR is created to enable detecting objects in extremely long point cloud sequences (hundreds of frames). In the MoDAR framework, a sequence is divided into several short sequences where objects are detected by a single-frame detector, tracked by a simple multi-object tracktor (e.g., [28]). Then, the data-driven motion forecasting model MultiPath++[29] predicts objects’ future poses based on objects’ trajectories established by the tracktor. Using prediction about objects’ future poses, detected objects in each short subsequence are propagated to the desired time step (e.g., the present). Next, each propagated object, which is represented as an up-right 3D bounding box (parameterized by the location of its center, size, and heading) with a confidence score and a class, is interpreted into a 3D point that takes the box’s center as its coordinate and the box’s size, heading, confidence score and class as features. These points interpreted from 3D bounding boxes are referred to as MoDAR points. They enable packing an entire subsequence into a small number of points, thus enabling an efficient fusion of extremely long point cloud sequences.
III-B2 V2X Collaboration using MoDAR Points
We draw the following similarity between single-vehicle object detection on point cloud sequences and V2X collaborative detection: these two tasks share the common challenge of finding an effective method for fusing information obtained from different perspectives caused by the motion of the ego vehicle in the case of point cloud sequences and the presence of other agents in the case of collaboration via V2X. Based on this observation, we use MoDAR points as the medium for conveying information among agents in the V2X network. Specifically, we interpret an object detected by agent $\mathcal{A}_{i}$ as a 3D bounding box $\mathbf{b}_{i,j}=[x,y,z,w,l,h,\theta,s,c]$ to a MoDAR point $\mathbf{m}_{i,j}$ by assigning
  • $[x,y,z]$ to the coordinate of $\mathbf{m}_{i,j}$
  • $[w,l,h,\theta,s,c]$ to $\mathbf{m}_{i,j}$’s features
The challenge in our V2X setting is that different agents detect objects at different rates, thus forcing the ego vehicle to utilize MoDAR points made by other agents at passed time steps. This timestamp mismatch results in a spatial misalignment between exchanged MoDAR points and ground truth dynamic objects, which can diminish the benefit of collaboration or even decrease the ego vehicle’s accuracy. This challenge is encountered in the context of single-vehicle detection on point cloud sequences as well because dynamic objects change their poses from one subsequence to another. [9] resolves this by predicting objects’ pose using a multi-object tracktor and the motion forecasting model MultiPath++.
While this is feasible in the V2X context, the implementation of add-on modules for future pose prediction does not align with our design target of minimal architecture. Instead, we use the scene flow to propagate MoDAR points from a passed timestep to the timestep queried by the ego vehicle. Since MoDAR points are virtual, their scene flow is not estimated directly from a point cloud sequence but is aggregated from the scene flow of points residing in the box they represent. To be concrete, let $\mathbf{m}_{i,j}$ be a MoDAR point representing a 3D bounding box $\mathrm{B}_{i,j}$ detected by agent $\mathcal{A}_{i}$ at timestep $t_{i}$. $\mathbf{p}_{i,h}$ and $\mathbf{f}_{i,h}$ respectively denote a real 3D point in the concatenation of agent $\mathcal{A}_{i}$’s point cloud sequence $\mathcal{S}_{i}$ and its predicted scene flow. The scene flow $\mathbf{f}^{\mathbf{m}}_{i,j}$ of $\mathbf{m}_{i,j}$ is computed by
$\mathbf{f}^{\mathbf{m}}_{i,j}=\mathrm{mean}\left\{\mathbf{f}_{i,h}~{}|~{}\mathbf{p}_{i,h}\in\mathrm{B}_{i,j}\right\}$ (4)
Once its scene flow is obtained, the MoDAR point $\mathbf{m}_{i,j}$ is propagated to the timestep $t$ queried by the ego vehicle as following
$[\hat{\mathbf{m}}_{i,j}]_{x,y,z}=[\mathbf{m}_{i,j}]_{x,y,z}+\frac{t-t_{i}}{|\mathcal{S}_{i}|}~{}\mathbf{f}^{\mathbf{m}}_{i,j}$ (5)
Here, $|\mathcal{S}_{i}|$ denotes the length of the point cloud sequence $\mathcal{S}_{i}$ measured in seconds. $[\cdot]_{x,y,z}$ is the operator that extracts 3D coordinate of a MoDAR point.
Finally, propagated MoDAR point is transformed from the agent $\mathcal{A}_{i}$’s pose $\mathcal{E}_{i}(t_{i})$ at time step $t_{i}$ to the ego vehicle pose $\mathcal{E}_{e}(t)$ at time step $t$ using the localization in the common global frame $\mathcal{G}$ of the two agents
[${}^{\mathcal{E}_{e}(t)}[\hat{\mathbf{m}}_{i,j}]_{x,y,z}=~{}^{\mathcal{G}}\mathbf{T}^{-1}_{\mathcal{E}_{e}(t)}~{}^{\mathcal{G}}\mathbf{T}_{\mathcal{E}_{i}(t_{i})}~{}[\hat{\mathbf{m}}_{i,j}]_{x,y,z}$ (6)
The concatenation between the set of MoDAR points received from other agents and the ego vehicle’s raw point cloud (resulting from concatenating its own point cloud sequence) is done straightforwardly by padding
  • points in the ego vehicle’s raw point cloud with null vectors representing features of MoDAR points, which are boxes’ size, heading, score, and class
  • MoDAR points with null vectors representing features of points in the ego vehicle’s raw point cloud which are points’ intensity and time-lag.
Once exchanged MoDAR points and the point cloud of the ego vehicle are merged, the result can be processed by the single-vehicle model developed in the Sec.III-A2 without changing its architecture.
IV Experiments and Results
IV-A Single-Vehicle Perception
IV-A1 Datasets and Metrics
As point clouds’ resolution exerts a large impact on the performance of LiDAR-based models, we test our model on three resolutions: 16, 32, and 64. While 32-channel and 64-channel point clouds are readily available in NuScenes dataset [2] and KITTI dataset [17] respectively, to the best of our knowledge, there are no publicly available datasets containing 16-channel point clouds. As a result, we synthesize a 16-channel dataset from the NuScenes dataset using the downsampling approach of [30].
The NuScenes dataset
is made of 850 20-second scenes split into 700 scenes for training and 150 scenes for validation. Each scene comprises data samples collected by a multimodal sensor suite including a 32-beam LiDAR operating at 20 Hz. In NuScenes’ convention, a keyframe is established once all sensors are in sync which happens every half a second. For each keyframe, objects are annotated as 3D bounding boxes. In addition, each object is assigned a unique ID that is kept consistent throughout a scene which enables us to generate the ground truth for the rectification transformation and scene flow. The downsampling of each point cloud in the NuScenes dataset to synthesize a 16-channel one is done by first identifying its points’ beam index via K-mean (K is 32 in the case of NuScenes) clustering on their azimuth coordinate, then assigning an equivalent beam index with respect to 16-channel LiDAR. More details can be found in [30]. Hereon, we refer to this synthesized dataset as NuScenes-16.
The KITTI dataset
’s 3D Object Detection partition contains 7481 and 7581 samples for training and testing. Each sample comprises sensory measurements collected by a 64-beam LiDAR operating at 10Hz and several cameras. A common practice when working with KITTI is to split the original training data into 3712 training samples and 3769 validation samples for experimental studies. A challenge we encounter when using KITTI is that its 3D Object Detection partition contains temporally disjointed samples, thus being not straightforward to obtain input (point cloud sequences) and ground truth (scene flow) for our model. We resolve this challenge by matching samples in the 3D Object Detection partition with KITTI’s raw sequences. Since there are a few raw sequences that do not have tracklet annotation, meaning objects are not tracked, we only retain data samples whose associated raw sequences have tracklets (i.e., trajectories of objects) in the training set and validation set.
Metrics
We use mean Average Precision (mAP) to measure the performance of our model on the 3D object detection task. A prediction is matched with the ground truth that is closest, measured by an affinity, to it. A match is considered valid if the affinity is below a predefined threshold. For each threshold, the average precision is obtained by integrating the recall-precision curve for recall and precision above 0.1. The mAP is the mean of the average precision of the threshold set. For NuScenes, the affinity is the Euclidean distance on the ground plane between centers of predictions and ground truth. This distance has four thresholds: 0.5, 1.0, 2.0, and 4.0. For KITTI, the affinity is the Intersection-over-Union (IoU) in the BEV plane. Unlike NuScenes, KITTI uses only one affinity threshold for each class which is 0.7 for cars and 0.5 for pedestrians.
The evaluation of scene flow prediction is based on the set of standard metrics proposed by [31] which includes End-Point Error (EPE), strict/ relaxed accuracy (AccS/ AccR), and outlier (ROutliers). The EPE is the Euclidean distance between the predicted scene flow and their ground truth average over the total number of points. The AccS/ AccR is the percentage of points having either EPE $<$ 0.05/ 0.10 meters or relative error $<$ 0.05/ 0.10. The ROutliers is the percentage of points whose EPE $>$ 0.30 meters and relative error $>$ 0.30.
IV-A2 Experiments and Results
The implementation of the single-vehicle perception model here follows the implementation made in our previous work [1] which uses the concatenation of 0.5-second point cloud sequences by EMC as the input. Since NuScenes and KITTI respectively obtain point clouds at 20 and 10 Hz, a 0.5-second sequence contains 10 and 5 point clouds. We use the same architecture, which uses PointPillar [22] as its backbone and CenterHead [4] as its RPN, for every experiment. In experiments on NuScenes, the detection range is limited to $[-51.2,51.2]$ on the XY plane and $[-5.0,3.0]$ along the Z axis. For experiments in KITTI, point clouds are first cropped using the left-color camera’s field of view as the annotations are only provided for objects that lie within the left-color camera’s image. The detection range on KITTI dataset is $[0,69.12]\times[-39.68,39.68]\times[-3,1]$ along X, Y, and Z axis. Our implementation is based on the OpenPCDet [32]. The details on the model’s hyperparameters can be found in our code release.
The evaluation of the Aligner++ on scene flow estimation task is done using the NuScenes dataset. The result shown in Tab.I indicates a significant improvement compared to our previous work [1] that reaches a state-of-art on accuracy-related metrics namely AccS, AccR, and ROutliers while maintaining low inference time. However, we have relatively high EPE this is because the evaluation is done for every point in the point cloud. This means the evaluation is carried out for both ground truth foreground points which have associated ground truth scene flow (can be nonzero or zero depending on dynamic) and ground truth background points which we assign the null vector as ground truth scene flow. Since the classification module of the Object Head inevitably makes false positive/ false negative foreground predictions, a number of background/ foreground points are predicted to have nonzero/ zero scene flow. Even though the portion of false predictions is small, as indicated by accuracy (AccS, AccR) and outlier metrics, the magnitude of their error is sufficiently large, due to the fact either the prediction or ground truth is zero, thus resulting in a large EPE.
TABLE I: performance of our model on scene flow metrics
Method EPE $\downarrow$ AccS $\uparrow$ AccR $\uparrow$ ROutliers $\downarrow$ Runtime $\downarrow$
FLOT [33] 1.216 3.0 10.3 63.9 2.01
NSFPrior [34] 0.707 19.3 37.8 32.0 63.46
PPWC-Net [35] 0.661 7.6 24.2 31.9 0.99
WsRSF [36] 0.539 17.9 37.4 22.9 1.46
PCAccumulation [37] 0.301 26.6 53.4 12.1 0.25
Aligner [1] 0.506 16.8 30.2 33.8 0.09
Aligner++ 0.616 46.4 66.6 6.8 0.10
The better scene flow estimation made by the results in a higher detection accuracy on the NuScenes dataset which can be seen in Tab.II. The integration of Aligner++, made of HD Map and distillation using the teacher-student framework, improves the mAP average over 10 classes of objects of a plain PointPillar by 6 points which is almost double the gain brought by Aligner (3.4 points). Interestingly, the distillation is responsible for most (5.6 out of 6 points) of the success of the Aligner++.
TABLE II: detection improvement due to our extension
PointPillar Aligner [1] HD Map Distilling mAP (avg
Oracle 10 classes)
39.0
42.4
39.4
45.0
The robustness of our Aligner++ is demonstrated by experiments on the KITTI and the synthetic NuScenes-16 dataset. As can be seen Tab.III and Tab.IV, the integration of Aligner++ consistently improves the accuracy of detecting objects in point cloud sequences. A common point between these two tables is that performance gain amounts to the Aligner++ is larger for vehicle-like classes (e.g., car, truck, or trailer). This is because the scene flow ground truth is generated using the rectification transformation (2) which is established based on the rigid motion assumption. This assumption holds for vehicle-like classes while being an oversimplification for pedestrians. As a result, the estimation of scene flow for vehicle-like classes is more accurate, thus higher detection accuracy. Another important result can be drawn from Tab.III and Tab.IV is that the effectiveness, in terms of detection accuracy, of using point cloud sequences over single point clouds persists on various LiDAR resolutions.
TABLE III: detection performance on kitti dataset
Class Model Sequence Length
1 2 3 4
Car PointPillar 66.54 70.44 69.66 69.88
+ Aligner++ / +3.91 +7.45 +6.3
Pedes- PointPillar 5.32 6.16 16.71 20.77
trian + Aligner++ / +4.82 +4.07 +3.6
TABLE IV: detection performance on nuscenes-16
Model PointPillar PointPillar
+ Aligner++
Seq Length 1 10 10
Car 56.14 79.84 80.81
Truck 26.99 48.64 49.39
Const. Vehicle 0.61 8.21 9.45
Bus 38.06 59.91 59.99
Trailer 11.05 26.15 29.03
Barrier 38.02 59.07 59.26
Motorcycle 12.70 41.14 41.72
Bicycle 0.11 15.22 18.40
Pedestrian 45.57 76.13 76.64
Traffic Cone 33.04 54.66 54.51
mAP (avg. 10 26.23 46.90 47.92
classes)
IV-B V2X Collaborative Perception - Experiments
IV-B1 Dataset and Metric
To evaluate our collaboration framework, we use the V2X-Sim 2.0 [18] which is made using CARLA and the traffic simulator SUMO. This dataset is made of 100 100-frame sequences of traffic taking place at intersections of three towns of CARLA which are Town 3, Town 4, and Town 5. Each sequence contains data samples recorded at 5 Hz. Each data sample comprises raw sensory measurements made by the ego vehicle, one to four CAVs, and an IRSU, which is placed at an elevated position that has a large minimally occluded field of view of the intersection. Every vehicle and the IRSU are equipped with a 32-channel LiDAR. All agents are in sync which results in the same timestamp of data that they collect.
The V2X-Sim 2.0 dataset provides object annotations for each data sample. The official training, validation, and testing split are made of temporally disjoint data samples from three towns chosen such that there is no overlap in terms of intersections. Since we need point cloud sequences as input to our models, we can’t use the official splits. Instead, we use sequences in Town 4 and Town 5 as the training set, and those in Town 3 as the validation set, thus ensuring there is no intersection overlap. This choice results in an 8900-data-sample training set and an 1100-data-sample validation set. Since this dataset follows the format of the NuScenes, we use NuScenes’ implementation of mean Average Precision (mAP) (details in Sec.IV-A1) to measure the performance of our framework and baselines.
IV-B2 Implementation, Experiments and Results
In the convention of the V2X-Sim dataset, the IRSU and the ego vehicle are respectively assigned the id of 0 and 1 while other CAVs get id ranging from 2 to 5. To test our collaboration method’s ability to handle asynchronous exchanged information, we set the time lag between the timestamp $t$ of the ego vehicle’s query and the timestamp $t_{i}$ of the detection $\mathcal{B}_{i}=\left\{\mathbf{b}_{i,j}\right\}$ made by agent $\mathcal{A}_{i}(i\in\{0,2,3,4,5\})$ to the time gap between two consecutive data sample of V2X-Sim which is $0.2$ seconds.
We benchmark our approach to collaborative perception against two extremes of the performance-bandwidth spectrum which are Late and Early collaboration. Late Collaboration achieves the minimal bandwidth usage by fusing the detection $\mathcal{B}_{1}=\left\{\mathbf{b}_{1,j}\right\}_{j=1}^{|\mathcal{B}_{1}|}$ made by the ego vehicle with the detection made by other agents $\left\{\mathcal{B}_{i}|i\in\{0,2,3,4,5\}\right\}$ using Non-Max Suppression. We evaluate this baseline under three settings including
  • asynchronous exchange where the gap between $t_{i}$ and $t$ is $0.2$ seconds as described above
  • asynchronous exchange with $\mathcal{B}_{i}$ propagated from $t_{i}$ to $t$ using scene flow by the procedure described in section III-B2
  • synchronous exchange where there is no gap between $t$ and $t_{i}$
On the other hand, Early Collaboration reaches high performance by exchanging the entire raw point cloud sequences $\mathcal{S}_{i}=\left\{\mathcal{P}^{t_{i}-K+1}_{i},\dots,\mathcal{P}^{t_{i}}_{i}\right\}$ collected by each agent $\mathcal{A}_{i}$ is the second baseline. To explore the upper bound of V2X collaborative perception, we only apply the synchronous exchange to this baseline.
In the implementation of our V2X collaboration framework, every agent uses the single-vehicle detection model that is developed in Sec.III-A2 with the absence of the teacher module. The architecture and hyper-parameters are kept unchanged as the experiments of the single-vehicle model in Sec.IV-A2. While previous works using the V2X-Sim dataset [14, 18] set the detection range to $[-32,32]$ meters along the X and Y axis, centered on the ego vehicle, we extend this range to $[-51.2,51.2]$ to better demonstrate the performance gain thanks to collaborative perception via V2X.
To verify the ability of enhancing single-vehicle perception of V2X collaboration, we evaluate our approach and baselines in the setting where ground truths are made of objects visible to the ego vehicle meaning their bounding boxes contain at least one point of the point cloud $\mathcal{P}_{e}^{t}$, which the ego vehicle obtained at the time step of query $t$. The result of this evaluation is summarized in Tab.V which shows that late collaboration in all three settings is not beneficial, as the best late collaboration is 4.42 mAP behind the single-agent perception. This is because True Positive (TP) detections made by other agents are counted as false positives if they are not visible to the ego vehicle. In addition, ill-localized but overly confident detections made by other agents can suppress good detections made by the ego vehicle, thus further reducing the overall performance. Remarkably, our collaboration approach using MoDAR points outperforms the early collaboration by 3.35 mAP while consuming the same amount of bandwidth as late collaboration. This reaffirms our design philosophy that a good mutli-agent collaborative perception framework can be made on the foundation of a good single-agent perception model and a simple collaboration method.
TABLE V: performance of collaboration methods on ground truth that are visible to the ego vehicle
Collaboration Method mAP Exchanged Info
Size (MByte)
None 65.71 0
Late - async 54.77 $0.01$
Late - async prop. 59.34 $0.01$
Late - sync 61.29 $0.01$
Mid - DiscoNet 69.98 -
Early 71.84 $33.95$
Ours - MoDAR 75.19 $0.01$
Beside enabling the ego vehicle more accurately detecting objects that are visible to itself, another great benefit of collaborative perception is to help the ego vehicle see the invisible objects - those do not contain any points of its point cloud $\mathcal{P}_{e}^{t}$, thus overcoming the challenge of occlusion and sparsity. To demonstrate this, we perform the relax the criterion of eligibility of ground truth such that they only need to contain at least one point emitted by any agents in the V2X network. The performance shown in Tab.VI indicates that on a larger set of ground truths, Late Collaboration does improve performance, compared to single-vehicle (no collaboration). The improvement is significant (at least $8.35$ mAP) even in the poorest setting where the set of exchanged detection $\left\{\mathcal{B}_{i}|i\in\{0,2,3,4,5\}\right\}$ is $0.2$ seconds behind the time of query, thus resulting a spatial misalignment between detected objects that are exchanged and their associated ground truths if the underlying objects are dynamic. This can be explained by a significant number of static and slow-moving objects present in an intersection whose past detections remain true positive at the present. When detections of dynamic objects are better accounted for as in the sync setting where agents exchange their detections at the same timestamp as the query, the performance is largely improved by $9.29$ mAP, compared to the async setting. However, this sync setting is unrealistic because different agents have different detection rates. Interestingly, propagating detections using scene flow as in the async prop. setting can reach $96.20\%$ the performance of the sync setting ($2.68$ mAP behind). This implies the effectiveness of scene flow estimated by our Aligner++. In this setting, our collaboration method using MoDAR no longer outperfroms the early collaboration baseline as some distant objects (with respect to the ego vehicle) are inevitably missed by connected agents, thus remaining invisible to the ego vehicles and being unable to be detected. However, our method still maintain favorable performance of $98.2\%$ of the early collaboration’s while attaining as low bandwidth usage as late collaboration.
TABLE VI: performance of collaboration methods on ground truth that are visible to at least one agents
Collaboration Method mAP
None 52.84
Late - async 61.19
Late - async prop. 67.80
Late - sync 70.48
Early 78.10
Ours - MoDAR 76.72
Ours - MoDAR 75.41
and Foreground

### v3

@@
 Abstract
-In this paper, we improve the single-vehicle 3D object detection models using LiDAR by extending their capacity to process point cloud sequences instead of individual point clouds.
-In this step, we extend our previous work [1] on rectification of the shadow effect in the concatenation of point clouds to boost the detection accuracy of multi-frame detection models.
-Our extension includes incorporating HD Map and distilling an Oracle model.
-Next, we further increase the performance of single-vehicle perception using multi-agent collaboration via Vehicle-to-everything (V2X) communication.
-We devise a simple yet effective collaboration method that achieves better bandwidth-performance tradeoffs than prior arts while minimizing changes made to single-vehicle detection models and assumptions on inter-agent synchronization.
-Experiments on the V2X-Sim dataset show that our collaboration method achieves 98% performance of the early collaboration while consuming the equivalent amount of bandwidth usage of late collaboration which is 0.03% of early collaboration.
-The code will be released at https://github.com/quan-dao/practical-collab-perception
+Occlusion is a major challenge for LiDAR-based object detection methods as it renders regions of interest unobservable to the ego vehicle.
+A proposed solution to this problem comes from collaborative perception via Vehicle-to-Everything (V2X) communication, which leverages a diverse perspective thanks to the presence of connected agents (vehicles and intelligent roadside units) at multiple locations to form a complete scene representation.
+The major challenge of V2X collaboration is the performance-bandwidth tradeoff which presents two questions (i) which information should be exchanged over the V2X network, and (ii) how the exchanged information is fused.
+The current state-of-the-art resolves to the mid-collaboration approach where Birds-Eye View (BEV) images of point clouds are communicated to enable a deep interaction among connected agents while reducing bandwidth consumption.
+While achieving strong performance, the real-world deployment of most mid-collaboration approaches are hindered by their overly complicated architectures and unrealistic assumptions about inter-agent synchronization.
+In this work, we devise a simple yet effective collaboration method based on exchanging the outputs from each agent that achieves a better bandwidth-performance tradeoff while minimising the required changes to the single-vehicle detection models.
+Moreover, we relax the assumptions used in existing state-of-the-art approaches about inter-agent synchronization to only require a common time reference among connected agents, which can be achieved in practice using GPS time.
+Experiments on the V2X-Sim dataset show that our collaboration method reaches 76.72 mean average precision which is 99% the performance of the early collaboration method while consuming as much bandwidth as the late collaboration (0.01 MB on average).
+The code will be released in https://github.com/quan-dao/practical-collab-perception.
 Index Terms:
-V2X Collaborative Perception, 3D Object Detection, Scene Flow.
+collaborative perception, V2X, 3D object detection, deep learning, LiDAR.
 I Introduction
-Object detection, which amounts to localizing objects as 3D bounding boxes, is a fundamental autonomous driving software stack module.
-It provides input to downstream tasks such as object tracking, motion forecast, and navigation.
-Due to the need for accurately localizing objects in 3D, state-of-the-art methods use LiDARs as the primary sensing modality.
-While significant advancements have been made, two challenges continuously plague LiDAR-based detection models: occlusion and sparsity, especially at long ranges.
-As occluded objects in one frame can be visible in others due to their motion and/or the ego vehicle’s motion, and sparse regions can become dense as the ego vehicle approaches, the utilization of point cloud sequences is an appealing candidate for boosting the boundary of 3D object detection.
-First introduced in [2] as a means for alleviating the extreme sparsity of point clouds obtained by a 32-beam LiDAR, the concatenation of point cloud sequences using ego vehicle pose; thus the name Ego Motion Compensation (EMC), is by far the most popular multi-frame method due to its simplicity.
-Yet, being highly effective [3, 4, 5, 6, 7].
-More recent works such as 3D-MAN [8], MoDAR [9], and MPPNet [10] are built upon this concept by dividing the input sequence into several subsequences whose point clouds are concatenated to produce preliminary detections.
-These preliminary detections are then fused to result in the final detections.
-The major drawback of EMC-concatenated point clouds is that their formulation only accounts for the motion of the ego vehicle while ignoring the motion of dynamic objects, thus their instantiating at multiple locations, which is referred to use the shadow effect [11], during the observation horizon as can be seen in Fig.1.
-This effect, which is essentially a misalignment in 3D space between objects’ points and their location, results in a misalignment in feature spaces such as on the bird-eye view image, limiting the performance gain brought by an increased number of points (thanks to point cloud concatenation) to stationary and slow-moving objects [8].
-In our previous work [1], we resolved this alignment explicitly by rectifying concatenated point clouds using estimated point-wise scene flow.
-The add-on module named Aligner, which we proposed, achieved competitive scene flow performance while substantially improving the detection accuracy, especially of large objects with a severe shadow effect.
-Figure 1: A dynamic car’s appearance on the concatenation of a 10-point-cloud sequence spanning 0.5 seconds. Points are colour-coded according to their time lag with respect to the present.
-Despite the power of point cloud sequences, the issue of occlusion and sparsity at long range persists as several regions remain invisible to the ego vehicle due to the geometry of the scene (e.g., a large object obstructs the ego vehicle’s field of view) or the traffic rule that limits the ego vehicle’s motion (e.g., the ego vehicle must stop at a red light).
-These issues can be safety critical when the ego vehicle navigates complex traffic, such as at intersections where a Waymo report indicates that a quarter of severe accidents involving their autonomous vehicles are due to occlusion [12].
-Therefore, in this work, we aim to boost further object detection made by a single-vehicle operating on point cloud sequences using the Vehicle-to-everything (V2X) collaborative perception.
-In the collaborative perception framework via V2X, Intelligent Road-Side Units (IRSU), which are equipped with advanced sensing systems (e.g., comparable to an autonomous vehicle) and located at elevated locations that maximize their field of view, and other vehicles can make measurements in regions that are inaccessible or partially occluded to others.
-Together they can come up with a complete representation of the environment.
-The enhanced perception capacity, thanks to V2X, comes with several new technical challenges; the most notorious among them is the performance-bandwidth tradeoff which revolves around two questions (i) which information to be broadcasted and (ii) how to fuse the exchanged information.
-This tradeoff establishes a design spectrum ranging from early collaboration, which exchanges raw sensory measurements - point clouds in the context of this paper, thus consuming the maximum bandwidth, to achieve the maximal performance to late collaboration that exchanges outputs (e.g., object detection as 3D bounding boxes) to achieves minimal bandwidth usage at the expense of limited performance.
-Naturally, researches on V2X collaboration frameworks [13, 14, 15, 16] are drawn toward the middle of this spectrum, thus the category’s name of mid-collaboration, where intermediate representations such as bird-eye view (BEV) images of agents’ surrounding environment are chosen for broadcasting in an attempt of balancing the two mutually excluding design targets.
+The 3D detection and localization of objects is a fundamental module of an autonomous driving software stack.
+It provides input to downstream tasks such as object tracking, motion prediction, and navigation.
+Due to the requirement of localizing objects accurately in 3D, state-of-the-art methods use LiDARs as a primary sensing modality.
+While significant advancements have been made, occlusion remains troublesome for LiDAR-based detection models, which makes regions of interest inaccessible to the ego vehicle.
+The issue of occlusion becomes particularly critical when navigating complex traffic scenarios, such as intersections, as autonomous vehicles are required to avoid collisions using a field of view that is severely reduced due to a large number of road users.
+The need for addressing occlusions at intersections is highlighted in a report published by Waymo [1] stating that two out of eight severe accidents involving their autonomous vehicles are due to occlusion in the intersections.
+Vehicle-to-Everything (V2X) collaborative perception is a promising solution to the occlusion challenge.
+Its core idea is to form a complete scene representation using measurements collected from multiple perspectives by leveraging the communication among multiple connected agents presenting at different locations.
+Connected agents can be either connected and automated vehicles (CAVs) or Intelligent Road-Side Units (IRSUs), which are advanced sensing systems strategically placed at elevated locations to have maximal coverage of regions where complex traffic scenarios take place.
+This enhanced perception capacity, thanks to V2X, comes with several new technical challenges; the most notorious among them being the performance-bandwidth tradeoff which presents two questions; (i) which information should be broadcast, and (ii) how the exchanged information should be fused.
+This tradeoff establishes a spectrum of solutions ranging from early to late collaboration.
+Raw measurements, which are point clouds in the context of this paper, are exchanged in the framework of early collaboration to maximally reduce the impact of occlusion, thus achieving the highest performance at the expense of spending a significant amount of bandwidth.
+On the other extreme, late collaboration exchanges high-level outputs (e.g., object detection as 3D bounding boxes) to minimize bandwidth usage, though this approach has been demonstrated to have reduced performance gains.
+In an attempt to balance the two mutually excluding design targets, research on V2X collaboration frameworks [2, 3, 4, 5] are drawn toward the middle of this spectrum, thus the category’s name of mid-collaboration, where intermediate representations such as BEV images of agents’ surrounding environment are chosen for broadcasting.
 While the motivation is just, most mid-collaboration methods require making substantial changes to the architecture of single-agent perception models to accommodate the fusion module where the combination of exchanged representations takes place.
-More importantly, most methods make strong assumptions about the synchronization among connected agents.
+More importantly, these methods make strong assumptions about data synchronization among connected agents.
 For example, DiscoNet [3] considers a perfectly synced setting where agents share the same clock, collect and process point clouds at the same rate and the transmission/ receiving of BEV images experience zero latency.
-V2VNet [13] and ViT-V2X [15] account for latency by postulating a global misalignment between exchanged BEV images and the BEV image of the ego vehicle caused by the movement of the ego vehicle between the time step it makes the requests for BEV images and the time step the exchanged BEV images are received.
-This means these works implicitly assume agents in the V2X network obtain point clouds synchronously.
-Taking a different approach to mid-collaboration, we aim for a more practical collaboration framework that emphasizes minimal
-  • Bandwidth consumption
-  • Changes made to single-agent models
-  • Inter-agent synchronization assumption
-We ground our design at minimal bandwidth consumption which can only be achieved by exchanging agents’ output which detected objects.
+V2VNet [2] and ViT-V2X [4] account for latency by postulating a global misalignment between exchanged BEV images and that of the ego vehicle.
+The cause of this misalignment is the movement of the ego vehicle between the time step it queries the V2X network and the time when the exchanged BEV images are received.
+This implicitly assumes that agents in the V2X network obtain point clouds synchronously.
+Another drawback of the current state-of-the-art of V2X collaborative perception is that only one point cloud per agent is used.
+Given that objects obscured in one frame may become visible in subsequent frames due to their movement or the motion of connected agents, and sparse regions might become dense as they draw nearer, harnessing sequences of point clouds is a compelling strategy to improve the performance of collaborative perception.
+In fact, the rich literature on multi-frame methods for single-vehicle object detection [6, 7, 8, 9, 10, 11] has confirmed the effectiveness of point cloud sequences as a simple concatenation of point clouds in a common frame can boost detection accuracy by approximate $30\%$ [6].
+Finally, mid-collaboration methods assume connected agents share a common type of detector which is impractical for real-world deployment.
+This assumption is critical because the domain gap between the BEV representation of point clouds made by different detectors severely affect their fusion [5].
+Prior work [5] resolves this challenge by introducing an additional module on top of those needed for the mid-collaboration to account for various differences in the BEV images made by SECOND [12] and PointPillar [13], thus further complicating the collaborative perception architecture.
+Aware of the aforementioned drawbacks of previous works on V2X collaborative perception, we seek a practical collaboration framework that emphasizes:
+  • Minimal bandwidth consumption
+  • Minimal changes made to single-agent models
+  • Minimal inter-agent synchronization assumptions
+  • Support heterogeneous detectors networks
+We aim our design at minimal bandwidth consumption, which can only be achieved by exchanging information about the objects detected by each agent.
 This design choice naturally satisfies the second design target as it dismisses the need for complex mid-representation fusion modules.
-We achieve the final target by only assuming that connected agents share a common time reference which is practically achievable using GPS time.
-The challenge posed by our relaxed inter-agent synchronization assumption is that information (i.e., detected objects) broadcasted by agents in the V2X network may never have the same timestamp as the query made by the ego vehicle.
+We achieve the third target by only assuming that connected agents share a common time reference which is practically achievable using GPS time.
+To reach the last target, we decide to perform the fusion of exchanged information in the input of the ego vehicle’s detector.
+The challenge posed by our relaxed inter-agent synchronization assumption is that information (i.e., detected objects) broadcast by agents in the V2X network may never have the same timestamp as the query made by the ego vehicle.
 In other words, the detections made by other agents that are available on the V2X network always have an older timestamp compared to the current timestamp of the ego vehicle.
-This timestamp mismatch results in a misalignment between exchanged detected objects and their associated ground truths (if the detections are true positive).
-We resolve this misalignment by predicting objects’ velocities simultaneously with their poses using the Aligner module developed by our previous work [1].
-Specifically, each detected object pools point-wise scene flow produced by the Aligner to compute its velocity, which is then used to propagate this object from one timestamp to another.
-As the scene flow plays a more critical role in the V2X collaborative perception, we first extend our Aligner by incorporating HD Maps into its framework and improving its supervision by distilling an Oracle model (details are in Sec.III-A2).
-The resulting module is dubbed Aligner++.
-Aware of the limitation of the late collaboration whose details are presented in Sec.II, we derive a novel collaboration strategy that fuses detections made by connected agents at a past timestamp with the current point cloud acquired by the ego vehicle using inspiration from the recent advance in multi-frame single-vehicle object detection [9].
+This timestamp mismatch results in a misalignment between exchanged detected objects and their associated ground truths (if the detections are true positives), thus risking the overall performance.
+Our solution to this issue lies in the information that prior works have neglected - point cloud sequences.
+We reason that objects detected in the past can be propagated to the present if their velocities are available.
+The prediction of objects’ velocities pertains to motion prediction [14, 9, 15] or object tracking [16] which can negate our minimal architecture changes design target.
+Since we assume agents produce predictions at least at the rate they acquire point clouds, we only need short-term (e.g., 0.1 seconds if point clouds are collected at 10Hz) velocity prediction, which can be computed using scene flow, rather than long-term prediction (e.g., 3 seconds [14]) offered by motion prediction or tracking.
+As a result, we use the plug-in module for scene flow estimation developed by our previous work [17].
+This choice effectively makes our V2X collaboration framework a multi-frame method, thus enabling each connected agent, as an individual, to enjoy a boost in detection accuracy as single-vehicle multi-frame methods do.
 This paper makes the following contributions:
-  • Extending our previous work [1] to further boosts single-vehicle object detection accuracy and achieves more accurate scene flow prediction
-  • Deriving a practical framework for V2X collaborative perception that consumes minimal bandwidth while achieving competitive performance compared to the early collaboration in the perfect synchronization setting
-  • Performing extensive evaluations on NuScenes [2], KITTI [17], and V2X-Sim [18] dataset to verify our method
-II Related Works
-A V2X cooperative perception framework comprises two main components connected autonomous vehicles (CAVs) and intelligent roadside units (IRSUs), both equipped with advanced sensing systems to perceive their environment.
-IRSUs are stationery installed at elevated positions to obtain minimally occluded fields of view.
-CAVs and IRSUs can exchange information to enhance their perception capacity (e.g., seeing beyond occlusion).
-In the context of this work, the sensing system of CAVs and IRSUs is restricted to a single LiDAR and the perception task focuses on 3D object detection.
-As described in the previous section, the main challenge of V2X cooperative perception is the performance-bandwidth tradeoff which establishes a design spectrum ranging from early to late collaboration.
+  • Deriving a practical framework for V2X collaborative perception that outperforms Early Collaboration while consuming as much bandwidth as Late Collaboration.
+In addition, our method does not make any assumptions about inter-agent synchronization except the existence of a common time reference, introduces minimal changes to the architecture of single-vehicle detectors, and supports heterogeneous detector networks.
+  • Demonstrate the benefit of point cloud sequences in V2X collaborative perception
+  • Extending our previous work [17] to further boost single-vehicle object detection accuracy and achieves more accurate scene flow prediction
+  • Performing extensive evaluations on NuScenes [6], KITTI [18], and V2X-Sim [19] datasets to validate the performance our method
+The remainder of this paper is organized as follows.
+Sec.II reviews the related works on V2X collaborative perception and highlights the gap to which we aim to contribute.
+Next, our methodology is explained in Sec.III.
+Sec.IV presents our experimental results.
+Finally, the conclusion and future perspective are drawn in Sec.V.
+II Related Work
+As described in the previous section, the main challenge of V2X cooperative perception is the performance-bandwidth tradeoff which establishes a solution spectrum ranging from early to late collaboration.
 In the early collaboration approach, as depicted in Fig.1, agents exchange their raw measurements - point clouds.
-At every timestep, the ego vehicle concatenates its own point cloud with those obtained by other CAVs and IRSUs to form the input for its perception module.
+At every timestep, the ego vehicle concatenates its own point cloud with those obtained by other agents to form the input for its perception model.
 This combined point cloud offers a comprehensive view of the scene with minimal occlusion and sparsity thanks to the diverse perspectives of connected agents, thus being regarded as the upper bound of the performance of the V2X collaborative perception [19, 20].
 However, due to the significant amount of bandwidth required to transmit raw point clouds (the order of 10 MB), the early collaboration strategy is not feasible for real-world deployments.
-Figure 2: Architecture of the early collaboration.
-On the other extreme of the performance-bandwidth tradeoff, the late collaboration approach focuses on minimizing bandwidth usage by exchanging only the agents’ predictions, specifically the detection results in the form of 3D bounding boxes.
-As illustrated in Fig.3, each agent independently conducts object detection on its own point cloud.
+Figure 1: Early Collaboration framework. Point clouds broadcast by connected agents are concatenated with that of the ego vehicle, forming the input to its detector.
+On the other extreme of the performance-bandwidth tradeoff, the late collaboration approach focuses on minimizing bandwidth usage by exchanging only the detection results from each agent in the form of 3D bounding boxes.
+As illustrated in Fig.2, each agent independently detects objects using its own point cloud.
 Subsequently, the agent merges its own predictions with those made by others to generate the final output.
@@
 In noisy environments where latency is a factor, the late collaboration strategy even underperforms single-vehicle perception [4].
-Figure 3: Architecture of the late collaboration.
+Figure 2: Late Collaboration framework. Objects detected by other agents are fused with those detected by the ego vehicle to form its final output.
 The mid-collaboration strategy aims to find a balance between performance and bandwidth consumption by exchanging intermediate scene representations generated by the backbone of the agents’ perception model.
 The motivation behind this approach is that the intermediate scene representation contains more contextual information compared to the final output (i.e., 3D bounding boxes), enabling deeper interaction among agents.
-Moreover, this representation is more compact than raw point clouds since it has been reduced in size through a series of convolution layers in the backbone.
-It can also be further compressed using an autoencoder to minimize bandwidth usage.
-While the idea is elegant, implementing the intermediate collaboration strategy requires a range of modules, shown in Figure 4, including compressor, decompressor, and representation fusion, among others, to match the performance of early collaboration.
+Moreover, this representation is more compact than raw point clouds since it has been reduced in size through a series of convolution layers in the backbone and can be further compressed using an autoencoder to minimize bandwidth usage.
+While the idea is elegant, implementing the intermediate collaboration strategy requires a range of modules, shown in Fig.3, including compressor, decompressor, and representation fusion, among others, to match the performance of early collaboration.
 The fusion, in particular, is quite intricate as it involves learnable collaboration graphs using techniques such as Graph Neural Networks [2, 3] or Transformers [4] to effectively fuse exchanged representations.
-Furthermore, dedicated modules are required to account for different practical challenges.
+Figure 3: Mid Collaboration framework. The ego vehicle fuses the BEV image of its point cloud with those broadcast by other agents to improve its performance.
+In addition, dedicated modules are required to account for different practical challenges.
 For example, [2, 4] use the Spatial Transformer [21] to resolve the global misalignment between the ego vehicle’s representation and others’ caused by the ego vehicle’s motion between when it makes the query and when it receives exchanged information.
-In [16], the Fused Axial Attention [21] is used to bridge the domain gap between representations made by different detection models (e.g., PointPillar [22] and VoxelNet[23]) used by different agents in the V2X network.
-More importantly, most mid-collaboration methods make strong assumptions about inter-agent synchronization which is either (i) perfect synchronization where exchanged representations always share the same timestamp [14] or (ii) synchronized point cloud acquisitions, meaning all agents obtain and process point clouds at the same rate and at the same time [13, 15].
+In [5], the Fused Axial Attention [22] is used to bridge the domain gap between representations made by different detection models (e.g., PointPillar [23] and SECOND[12]) used by different agents in the V2X network.
+Finally, most mid-collaboration methods make strong assumptions about inter-agent synchronization which is either (i) perfect synchronization where exchanged representations always share the same timestamp [3] or (ii) synchronized point cloud acquisitions, meaning all agents obtain and process point clouds at the same rate and at the same time [2, 4].
 Because of these complexities, the real-world deployment of intermediate fusion remains challenging.
-Figure 4: V2X Cooperative Perception via the Mid-fusion architecture.
+III Methodology
 In this work, we aim to resolve the aforementioned complexities of mid-collaboration to obtain a practical framework for V2X collaborative perception.
-Our design is grounded in minimal bandwidth consumption as this is nonnegotiable in real-world deployment.
-To achieve this design goal, we choose object detection, in the form of 3D bounding boxes, as the information to be exchanged, which is similar to the late collaboration strategy.
+Our design is grounded in minimal bandwidth consumption as this is essential for real-world deployment at scale.
+To achieve this goal we choose object detection, in the form of 3D bounding boxes, as the information to be exchanged which is similar to the late collaboration strategy.
 We further relax the assumption on inter-agent synchronization to agents sharing a common time reference (e.g., GPS time) and acknowledge that agents produce detections at different rates.
 As a result, exchanged detections always have older timestamps compared to the timestamp of the query made by the ego vehicle, thus risking a spatial misalignment between exchanged detections and their associated ground truth (if detections are true positive).
-We resolve this misalignment by predicting objects’ velocity simultaneously with their locations by pooling point-wise scene flow which can be produced by integrating our Aligner++ module to any off-the-shelf BEV-based object detectors.
-Finally, we avoid the inferior performance of late collaboration by devising a new collaboration strategy that fuses exchanged detections with the ego vehicle’s raw point cloud for subsequent processing using its detection model rather than fusing exchanged detections with detections made by the ego vehicle.
-An illustration of our approach is presented in Fig.5.
-The innovation of our collaboration strategy lies in recognizing the similarity between object detection using point cloud sequences and collaborative detection.
+We resolve this misalignment by predicting objects’ velocity simultaneously with their locations by pooling point-wise scene flow which can be produced by integrating our Aligner module [17] to any BEV-based object detectors.
+Finally, we avoid the inferior performance of late collaboration by devising a new collaboration strategy that fuses exchanged detections with the ego vehicle’s raw point cloud for subsequent processing by its detection model.
+The resulting framework is illustrated in Fig.4.
+We name our method late-early collaboration as connected agents broadcast their outputs, which is the signature of late collaboration, while the ego vehicle fuse received information with its point cloud, which is the signature of early collaboration.
+Figure 4: Our late-early collaboration framework for V2X Cooperative Perception.
+In which, objects detected by each connected agent at a past time that is closest to the present are broadcast.
+Exchanged detected objects are propagated to the present using their velocity computed by pooling point-wise scene flow, then fused with the point cloud which the ego vehicle collected at the present to enhance its perception.
+The innovation of our collaboration strategy lies in our recognition of the similarity between object detection using point cloud sequences and collaborative detection.
 In both cases, there is a need to fuse information obtained from diverse perspectives.
 Point cloud sequences involve capturing the motion of the ego vehicle, which results in varying viewpoints, while collaborative detection entails incorporating insights from other agents present in the environment.
-By drawing this parallel, we leverage the shared principle of fusing information from multiple perspectives to enhance the accuracy and robustness of both object detection approaches
-Figure 5: Our V2X Cooperative Perception.
-III Methodology
-This section first presents the extension we introduce to our previous work [1] to improve the scene flow estimation as well as further boost the performance gain when integrating our module to single-vehicle object detection models.
+By drawing this parallel, we leverage the shared principle of fusing information from multiple perspectives to enhance the accuracy and robustness of both object detection approaches.
+Specifically, we utilize the latest advancement in the multi-frame object detection literature, called MoDAR [24], which interprets previously detected objects to 3D points with additional features made of object sizes, heading, confident score, and predicted class.
+These points are propagated to the present and merged with the point cloud obtained at the present to form the input of any off-the-shelf detectors.
+Since the scene flow plays a prominent role in our V2X collaboration framework by propagating past detections to the present time, this section first recalls our previous work [17] and presents the extension we make to improve the accuracy of its scene flow prediction.
 Next, our strategy for V2X collaborative perception that enables using asynchronous exchanged information is derived.
-III-A Improving Single-Vehicle Perception
-As mentioned in Sec.I, the accuracy of single-vehicle detection can be significantly improved by using the concatenation point cloud sequences as inputs.
-However, such concatenation is done by the Ego Motion Compensation method which only accounts for the motion of the ego vehicle resulting in the shadow effect (Fig.1) in the concatenated point cloud.
+III-A Learning Scene Flow
+As mentioned in Sec.I, the accuracy of single-vehicle detection can be significantly improved by using concatenated point cloud sequences as inputs.
+However, such concatenation is done by the Ego Motion Compensation (EMC) method which only accounts for the motion of the ego vehicle resulting in the shadow effect, as can be seen in Fig.5, in the concatenated point cloud.
 This effect, which is a misalignment in the 3D space between objects’ points and their locations, leads to a misalignment in feature spaces (e.g., BEV representations), thus degrading detection accuracy for dynamic objects.
 In this section, we recall the approach we use in [17] to handle this effect and present the extensions we make to it.
-III-A1 Shadow Effect: Cause and Rectification
+Figure 5: A dynamic car’s appearance on the concatenation of a 10-point-cloud sequence spanning 0.5 seconds. Points are color-coded according to their time lag with respect to the present.
+III-A1 The Cause of The Shadow Effect
 Let $\mathcal{P}^{t}=\left\{\mathbf{p}_{j}^{t}=[x,y,z,r,\Delta]~{}|~{}j=1,\dots,N\right\}$ denote a point cloud collected by the ego vehicle at time step $t$.
-Here, each point $\mathbf{p}_{j}$ has two features namely reflectant $r$ and time-lag $\Delta$ with respect to a predefined time step.
+Here, each point $\mathbf{p}_{j}$ has two features namely intensity $r$ and time-lag $\Delta$ with respect to a predefined time step.
 The point cloud is expressed in the ego vehicle frame $\mathcal{E}(t)$ measured with respect to a global frame $\mathcal{G}$.
-To concatenate or a point cloud sequence $\mathcal{S}=\left\{\mathcal{P}^{t-K},\mathcal{P}^{t-K+1},\dots,\mathcal{P}^{t}\right\}$ of length $K+1$, each point of point cloud $\mathbf{P}^{t-\Delta t}$ is transformed to the global frame using the ego vehicle pose at the same time step as in (1), hence the method’s name Ego Motion Compensation (EMC).
+To concatenate a point cloud sequence $\mathcal{S}=\left\{\mathcal{P}^{t-K},\mathcal{P}^{t-K+1},\dots,\mathcal{P}^{t}\right\}$ of length $K+1$, each point of point cloud $\mathbf{P}^{t-\Delta t}$ is transformed to the global frame using the ego vehicle pose at the same time step as in (1), hence the method’s name Ego Motion Compensation.
 $^{\mathcal{G}}\mathbf{p}_{i}^{t-\Delta t}=~{}^{\mathcal{G}}\mathbf{T}_{\mathcal{E}(t-\Delta t)}\cdot~{}\mathbf{p}_{i}^{t-\Delta t}$ (1)
 Here, ${}^{d}\mathbf{T}_{s}\in\mathbf{SE(3)}$ represents the rigid transformation that maps points in frame $s$ to frame $d$.
-EMC undoes the motion of the ego vehicle; however, it fails to account for objects’ motion.
+EMC undoes the motion of the ego vehicle without accounting for objects’ motion.
 As a result, the appearance of a dynamic object in the concatenated point cloud comprises several instances, each corresponding to the object’s poses in the global frame at a particular time step, as in Fig.5.
@@
 III-A2 Learning to Rectify the Shadow Effect
-In [1], we propose a drop-in module, named Aligner, for 3D object detection models that create a Bird-Eye View (BEV) representation, $\mathbf{B}\in\mathbb{R}^{C\times H\times W}$, of the scene as an intermediate output.
-Our module first computes the feature of $\mathbf{f}_{i}\in\mathbb{R}^{C}$ of a point $\mathbf{p}_{i}$ by bilinear interpolating on $\mathbf{B}$ using its projection to the BEV.
+In [17], we propose a drop-in module, named Aligner, for 3D object detection models that creates a BEV representation, $\mathbf{B}\in\mathbb{R}^{C\times H\times W}$, of the scene as an intermediate output.
+Our module first computes the feature of $\mathbf{f}_{i}\in\mathbb{R}^{C}$ of a point $\mathbf{p}_{i}$ by the bilinear interpolation of $\mathbf{B}$ using its projection to the BEV.
 Each point feature is then decoded into the point’s scene flow $\mathbf{d}_{i}\in\mathbb{R}^{3}$ representing how much this point needs to offset to rectify the shadow effect.
@@
 The complete pipeline is summarized in Fig.6.
-Figure 6: Aligner
+Figure 6: Aligner’s overall pipeline.
+Point-wise scene flow is predicted by Point Head based on point-wise features which are interpolated from the BEV image of the input point cloud.
+The input point cloud is then rectified by adding the corresponding scene flow to the coordinate of each LiDAR point.
+The scattering of point-wise features to the BEV using their rectified coordinate results in the rectified BEV image which is shadow-effect-free but sparse.
+The original BEV image which suffers from the shadow effect and the rectified BEV image are fused to minimize the shadow effect and sparsity.
+The resulting fused BEV image is finally fed to the RPN to obtain object detection.
 The learning target $\mathbf{d}^{*}$ for the scene flow $\mathbf{d}$ of a point $\mathbf{p}^{t-\Delta t}$, which has the timestamp $(t-\Delta t)$ (i.e., collected at time step $t-\Delta t$), is the difference between its location and its rectified location computed by (2).
@@
 The learning target is defined according to the framework which the RPN follows (e.g., anchor-based [12] or center-based [8]).
-To improve the accuracy of [1] in estimating scene flow, thus ultimately improving object detection, we introduce two extensions (i) incorporating HD Map and (ii) distilling an Oracle model.
-Previous works in integrating HD Maps into 3D object detection models ([26, 27]) opt for a mid-fusion approach that concatenates rasterized maps with backbone-made BEV representations.
+III-A3 Aligner++
+To improve the accuracy of [17] in estimating scene flow, thus ultimately improving object detection, we introduce two extensions (i) incorporating HD Map and (ii) distilling a model trained on the concatenation of point cloud sequences by the ground truth trajectories.
+Previous works in integrating HD Maps into 3D object detection models [26, 27] opt for a mid-fusion approach that concatenates rasterized maps with backbone-made BEV representations.
 However, this approach is incompatible with copy-paste data augmentation [12], which randomly samples ground truth objects from a database and pastes them to each point cloud used for training, as pasted objects do not necessarily adhere to the semantics and geometry of the map (e.g., cars appear inside a building).
@@
 As a result, HD Maps are made compatible with copy-paste data augmentation.
-Figure 7: Extracting map features for points.
-Figure 8: Alinger++ ’s Teacher-student framework
+Figure 7: Extraction of map features. LiDAR points are projected to the BEV. Their map feature is obtained by nearest-neighbor interpolation on the rasterized HD Map.
+Figure 8: Alinger++ ’s teacher-student framework.
+A model, pre-trained on point clouds concatenated using object ground truth trajectories, plays the role of the teacher.
+During training, the student model, which operates on point clouds concatenated using ego vehicle localization, strives to emulate the BEV images made by the teacher by minimizing a distillation loss.
 To improve the quality of the fused BEV representation, $\mathbf{B}^{f}$ in Fig.6, in terms of minimizing the feature misalignment caused by the shadow effect and reducing the sparsity, we use the teacher-student framework shown in Fig.8.
 In detail, we use ground truth objects’ trajectories to rectify the concatenation of point cloud sequences and use the shadow-effect-free results, illustrated in Fig.9, to train a teacher model.
-After the teacher converged, its BEV representation $\mathbf{B}^{t}$ is used to guide the student’s fused BEV representation $\mathbf{B}^{f}$ by optimizing the students’ weight so that the difference measured by the $\mathit{L}_{2}$ loss between these two representations is minimized.
+After the teacher converges, its BEV representation $\mathbf{B}^{t}$ is used to guide the student’s fused BEV representation $\mathbf{B}^{f}$ by optimizing the students’ weight so that the difference measured by the $\mathit{L}_{2}$ loss between these two representations is minimized.
 Figure 9: Comparison between concatenation of a point cloud sequence (a) and its rectification using objects’ ground truth trajectories (b).
-Points are color-coded according to their time-lag with respect to the present.
+Points are color-coded according to their time lag with respect to the present.
 The color bar is shown in (c).
@@
 At a time step $t_{i}$, agent $\mathcal{A}_{i}$ uses a $K$-point-cloud sequence $\mathcal{S}_{i}=\left\{\mathcal{P}^{t_{i}-K+1}_{i},\dots,\mathcal{P}^{t_{i}}_{i}\right\}$ as an input to its detection model.
-An object $b_{i,j}$ detected by agent $\mathcal{A}_{i}$ is parameterized by a nine-vector $[x,y,z,w,l,h,\theta,s,c]$.
+An object $\mathbf{b}_{i,j}$ detected by agent $\mathcal{A}_{i}$ is parameterized by a nine-vector $[x,y,z,w,l,h,\theta,s,c]$.
 The first seven numbers localize the object by its center location $[x,y,z]$, size $[w,l,h]$, heading direction $\theta$.
-The last two numbers $s$ and $c$ respectively denote confidence score $s$ and the predicted class $c$.
+The last two numbers, $s$ and $c$, respectively denote confidence score $s$ and the predicted class $c$.
 Upon receiving a query having timestamp $t$, agent $\mathcal{A}_{i}$ will communicate its detection $\mathcal{B}_{i}^{t_{i}}=\left\{\mathbf{b}_{i,j}\right\}_{j=1}^{M_{i}}$ and metadata produced timestamp $t_{i}$ that is prior to and closest to $t$.
@@
 MoDAR is created to enable detecting objects in extremely long point cloud sequences (hundreds of frames).
-In the MoDAR framework, a sequence is divided into several short sequences where objects are detected by a single-frame detector, tracked by a simple multi-object tracktor (e.g., [28]).
-Then, the data-driven motion forecasting model MultiPath++[29] predicts objects’ future poses based on objects’ trajectories established by the tracktor.
+In the MoDAR framework, a sequence is divided into several short sequences where objects are detected by a single-frame detector, tracked by a simple multi-object tracker (e.g., [16]).
+Then, the data-driven motion forecasting model MultiPath++[15] predicts objects’ future poses based on objects’ trajectories established by the tracker.
 Using prediction about objects’ future poses, detected objects in each short subsequence are propagated to the desired time step (e.g., the present).
-Next, each propagated object, which is represented as an up-right 3D bounding box (parameterized by the location of its center, size, and heading) with a confidence score and a class, is interpreted into a 3D point that takes the box’s center as its coordinate and the box’s size, heading, confidence score and class as features.
-These points interpreted from 3D bounding boxes are referred to as MoDAR points.
+Next, each propagated object, which is represented as an up-right 3D bounding box (parameterized by the location of its center, size, and heading) with a confidence score and a class, is converted into a 3D point that takes the box’s center as its coordinate and the box’s size, heading, confidence score and class as features.
+These points are referred to as MoDAR points.
 They enable packing an entire subsequence into a small number of points, thus enabling an efficient fusion of extremely long point cloud sequences.
@@
 This challenge is encountered in the context of single-vehicle detection on point cloud sequences as well because dynamic objects change their poses from one subsequence to another.
-[9] resolves this by predicting objects’ pose using a multi-object tracktor and the motion forecasting model MultiPath++.
-While this is feasible in the V2X context, the implementation of add-on modules for future pose prediction does not align with our design target of minimal architecture.
-Instead, we use the scene flow to propagate MoDAR points from a passed timestep to the timestep queried by the ego vehicle.
+[24] resolves this by predicting objects’ pose using a multi-object tracker and the motion forecasting model MultiPath++.
+While this is feasible in the V2X context, the implementation of those modules for future pose prediction does not align with our design target of minimal architecture.
+Instead, we use the scene flow to propagate MoDAR points from a prior timestep to the timestep queried by the ego vehicle.
 Since MoDAR points are virtual, their scene flow is not estimated directly from a point cloud sequence but is aggregated from the scene flow of points residing in the box they represent.
-To be concrete, let $\mathbf{m}_{i,j}$ be a MoDAR point representing a 3D bounding box $\mathrm{B}_{i,j}$ detected by agent $\mathcal{A}_{i}$ at timestep $t_{i}$.
+Specifically, let $\mathbf{m}_{i,j}$ be a MoDAR point representing a 3D bounding box $\mathbf{b}_{i,j}$ detected by agent $\mathcal{A}_{i}$ at timestep $t_{i}$.
 $\mathbf{p}_{i,h}$ and $\mathbf{f}_{i,h}$ respectively denote a real 3D point in the concatenation of agent $\mathcal{A}_{i}$’s point cloud sequence $\mathcal{S}_{i}$ and its predicted scene flow.
 The scene flow $\mathbf{f}^{\mathbf{m}}_{i,j}$ of $\mathbf{m}_{i,j}$ is computed by
-$\mathbf{f}^{\mathbf{m}}_{i,j}=\mathrm{mean}\left\{\mathbf{f}_{i,h}~{}|~{}\mathbf{p}_{i,h}\in\mathrm{B}_{i,j}\right\}$ (4)
+$\mathbf{f}^{\mathbf{m}}_{i,j}=\mathrm{mean}\left\{\mathbf{f}_{i,h}~{}|~{}\mathbf{p}_{i,h}\in\mathbf{b}_{i,j}\right\}$ (4)
 Once its scene flow is obtained, the MoDAR point $\mathbf{m}_{i,j}$ is propagated to the timestep $t$ queried by the ego vehicle as following
@@
 IV Experiments and Results
+This section first shows the results of the improvements to single vehicle perception with scene flow using real-world datasets, then demonstrates the collaborative approach using a simulated V2X dataset.
 IV-A Single-Vehicle Perception
 IV-A1 Datasets and Metrics
-As point clouds’ resolution exerts a large impact on the performance of LiDAR-based models, we test our model on three resolutions: 16, 32, and 64.
+As point clouds’ resolution has a large impact on the performance of LiDAR-based models, we test our model on three resolutions: 16, 32, and 64.
 While 32-channel and 64-channel point clouds are readily available in NuScenes dataset [6] and KITTI dataset [18] respectively, to the best of our knowledge, there are no publicly available datasets containing 16-channel point clouds.
@@
 The KITTI dataset
-’s 3D Object Detection partition contains 7481 and 7581 samples for training and testing.
-Each sample comprises sensory measurements collected by a 64-beam LiDAR operating at 10Hz and several cameras.
+has a 3D Object Detection partition containing 7481 and 7581 samples for training and testing.
+Each sample comprises sensor measurements collected by a 64-beam LiDAR operating at 10Hz and several cameras.
 A common practice when working with KITTI is to split the original training data into 3712 training samples and 3769 validation samples for experimental studies.
@@
 We use mean Average Precision (mAP) to measure the performance of our model on the 3D object detection task.
-A prediction is matched with the ground truth that is closest, measured by an affinity, to it.
+A prediction is matched with the closest ground truth, measured by an affinity.
 A match is considered valid if the affinity is below a predefined threshold.
@@
 For KITTI, the affinity is the Intersection-over-Union (IoU) in the BEV plane.
-Unlike NuScenes, KITTI uses only one affinity threshold for each class which is 0.7 for cars and 0.5 for pedestrians.
+Unlike NuScenes, KITTI uses only one affinity threshold for each class, which is 0.7 for cars and 0.5 for pedestrians.
 The evaluation of scene flow prediction is based on the set of standard metrics proposed by [29] which includes End-Point Error (EPE), strict/ relaxed accuracy (AccS/ AccR), and outlier (ROutliers).
@@
 For experiments in KITTI, point clouds are first cropped using the left-color camera’s field of view as the annotations are only provided for objects that lie within the left-color camera’s image.
-The detection range on KITTI dataset is $[0,69.12]\times[-39.68,39.68]\times[-3,1]$ along X, Y, and Z axis.
+The detection range on KITTI dataset is $[0,69.12]\times[-39.68,39.68]\times[-3,1]$ along the X, Y, and Z axis.
 Our implementation is based on the OpenPCDet [30].
 The details on the model’s hyperparameters can be found in our code release.
-The evaluation of the Aligner++ on scene flow estimation task is done using the NuScenes dataset.
-The result shown in Tab.I indicates a significant improvement compared to our previous work [1] that reaches a state-of-art on accuracy-related metrics namely AccS, AccR, and ROutliers while maintaining low inference time.
-However, we have relatively high EPE this is because the evaluation is done for every point in the point cloud.
+Scene Flow Evaluation
+The evaluation of the Aligner++ on scene flow estimation is done using the NuScenes dataset.
+The result shown in Tab.I indicates a significant improvement compared to our previous work [17] that reaches state-of-the-art on accuracy-related metrics namely AccS, AccR, and ROutliers while maintaining low inference time.
+The runtime of our methods, Aligner and Aligner++, is measured on an NVIDIA A6000 GPU, while the runtime of five baselines is measured on an NVIDIA RTX 3090 GPU (as reported by [31]) having the same compute capability.
+The Aligner++ has relatively high EPE; this is because the evaluation is done for every point in the point cloud.
 This means the evaluation is carried out for both ground truth foreground points which have associated ground truth scene flow (can be nonzero or zero depending on dynamic) and ground truth background points which we assign the null vector as ground truth scene flow.
@@
 PCAccumulation [31] 0.301 26.6 53.4 12.1 0.25
-Aligner [1] 0.506 16.8 30.2 33.8 0.09
-Aligner++ 0.616 46.4 66.6 6.8 0.10
+PointPillar + 0.506 16.8 30.2 33.8 0.09
+Aligner [17]
+PointPillar + 0.616 46.4 66.6 6.8 0.10
+Aligner++
+Object Detection Evaluation
 The better scene flow estimation made by the results in a higher detection accuracy on the NuScenes dataset which can be seen in Tab.II.
-The integration of Aligner++, made of HD Map and distillation using the teacher-student framework, improves the mAP average over 10 classes of objects of a plain PointPillar by 6 points which is almost double the gain brought by Aligner (3.4 points).
+The integration of Aligner++, made using the HD Map and through distillation using the teacher-student framework, improves the mAP average over 10 classes of objects of a plain PointPillar by 6 points which is almost double the gain brought by Aligner (3.4 points).
 Interestingly, the distillation is responsible for most (5.6 out of 6 points) of the success of the Aligner++.
@@
 PointPillar Aligner [17] HD Map Distilling mAP (avg
-Oracle 10 classes)
+Teacher 10 classes)
 39.0
 42.4
-39.4
+42.8
 45.0
@@
 As can be seen Tab.III and Tab.IV, the integration of Aligner++ consistently improves the accuracy of detecting objects in point cloud sequences.
-A common point between these two tables is that performance gain amounts to the Aligner++ is larger for vehicle-like classes (e.g., car, truck, or trailer).
+A common point between these two tables is that performance gain from Aligner++ is larger for vehicle-like classes (e.g., car, truck, or trailer).
 This is because the scene flow ground truth is generated using the rectification transformation (2) which is established based on the rigid motion assumption.
 This assumption holds for vehicle-like classes while being an oversimplification for pedestrians.
-As a result, the estimation of scene flow for vehicle-like classes is more accurate, thus higher detection accuracy.
-Another important result can be drawn from Tab.III and Tab.IV is that the effectiveness, in terms of detection accuracy, of using point cloud sequences over single point clouds persists on various LiDAR resolutions.
+As a result, the estimation of scene flow for vehicle-like classes is more accurate, thus resulting in higher detection accuracy.
+Another important result that can be drawn from Tab.III and Tab.IV is the effectiveness, in terms of detection accuracy, of using point cloud sequences over single point clouds persists across various LiDAR resolutions.
 TABLE III: detection performance on kitti dataset
@@
 IV-B1 Dataset and Metric
-To evaluate our collaboration framework, we use the V2X-Sim 2.0 [18] which is made using CARLA and the traffic simulator SUMO.
+To evaluate our collaboration framework, we use the V2X-Sim 2.0 [19] which is made using CARLA [36] and the traffic simulator SUMO [37].
 This dataset is made of 100 100-frame sequences of traffic taking place at intersections of three towns of CARLA which are Town 3, Town 4, and Town 5.
@@
 IV-B2 Implementation, Experiments and Results
-In the convention of the V2X-Sim dataset, the IRSU and the ego vehicle are respectively assigned the id of 0 and 1 while other CAVs get id ranging from 2 to 5.
+In the convention of the V2X-Sim dataset, the IRSU and the ego vehicle are respectively assigned the identity of 0 and 1 while other CAVs get identities ranging from 2 to 5.
 To test our collaboration method’s ability to handle asynchronous exchanged information, we set the time lag between the timestamp $t$ of the ego vehicle’s query and the timestamp $t_{i}$ of the detection $\mathcal{B}_{i}=\left\{\mathbf{b}_{i,j}\right\}$ made by agent $\mathcal{A}_{i}(i\in\{0,2,3,4,5\})$ to the time gap between two consecutive data sample of V2X-Sim which is $0.2$ seconds.
+In the implementation of our V2X collaboration framework, every agent uses the single-vehicle detection model that is developed in Sec.III-A2 with the absence of the teacher module.
+The architecture and hyper-parameters are kept unchanged as the experiments of the single-vehicle model in Sec.IV-A2.
+While previous works using the V2X-Sim dataset [3, 19] set the detection range to $[-32,32]$ meters along the X and Y axis, centered on the ego vehicle, we extend this range to $[-51.2,51.2]$ to better demonstrate the performance gain thanks to collaborative perception via V2X.
+Enhance Detection of Visible Objects
 We benchmark our approach to collaborative perception against two extremes of the performance-bandwidth spectrum which are Late and Early Collaboration.
@@
 On the other hand, Early Collaboration reaches high performance by exchanging the entire raw point cloud sequences $\mathcal{S}_{i}=\left\{\mathcal{P}^{t_{i}-K+1}_{i},\dots,\mathcal{P}^{t_{i}}_{i}\right\}$ collected by each agent $\mathcal{A}_{i}$ is the second baseline.
-To explore the upper bound of V2X collaborative perception, we only apply the synchronous exchange to this baseline.
-In the implementation of our V2X collaboration framework, every agent uses the single-vehicle detection model that is developed in Sec.III-A2 with the absence of the teacher module.
-The architecture and hyper-parameters are kept unchanged as the experiments of the single-vehicle model in Sec.IV-A2.
-While previous works using the V2X-Sim dataset [14, 18] set the detection range to $[-32,32]$ meters along the X and Y axis, centered on the ego vehicle, we extend this range to $[-51.2,51.2]$ to better demonstrate the performance gain thanks to collaborative perception via V2X.
-To verify the ability of enhancing single-vehicle perception of V2X collaboration, we evaluate our approach and baselines in the setting where ground truths are made of objects visible to the ego vehicle meaning their bounding boxes contain at least one point of the point cloud $\mathcal{P}_{e}^{t}$, which the ego vehicle obtained at the time step of query $t$.
-The result of this evaluation is summarized in Tab.V which shows that late collaboration in all three settings is not beneficial, as the best late collaboration is 4.42 mAP behind the single-agent perception.
+TABLE V: performance of collaboration methods on ground truth that are visible to the ego vehicle
+Collaboration Method mAP
+None 65.71
+Late - async 54.77
+Late - async prop. 59.34
+Late - sync 61.29
+Early 71.84
+Ours 75.19
+To verify the ability to enhance the single-vehicle perception of V2X collaboration, we evaluate our approach and baselines in the setting where ground truths are made of objects visible to the ego vehicle.
+This means eligible ground truth must contain at least one point of the point cloud $\mathcal{P}_{e}^{t}$, which the ego vehicle obtained at the time step of query $t$.
+The result of this evaluation is summarized in Tab.V which shows that Late Collaboration in all three settings is not beneficial, as the best Late Collaboration is only $93.3\%$ (4.42 mAP behind) of the single-agent perception.
 This is because True Positive (TP) detections made by other agents are counted as false positives if they are not visible to the ego vehicle.
 In addition, ill-localized but overly confident detections made by other agents can suppress good detections made by the ego vehicle, thus further reducing the overall performance.
-Remarkably, our collaboration approach using MoDAR points outperforms the early collaboration by 3.35 mAP while consuming the same amount of bandwidth as late collaboration.
-This reaffirms our design philosophy that a good mutli-agent collaborative perception framework can be made on the foundation of a good single-agent perception model and a simple collaboration method.
-TABLE V: performance of collaboration methods on ground truth that are visible to the ego vehicle
-Collaboration Method mAP Exchanged Info
-Size (MByte)
-None 65.71 0
-Late - async 54.77 $0.01$
-Late - async prop. 59.34 $0.01$
-Late - sync 61.29 $0.01$
-Mid - DiscoNet 69.98 -
-Early 71.84 $33.95$
-Ours - MoDAR 75.19 $0.01$
-Beside enabling the ego vehicle more accurately detecting objects that are visible to itself, another great benefit of collaborative perception is to help the ego vehicle see the invisible objects - those do not contain any points of its point cloud $\mathcal{P}_{e}^{t}$, thus overcoming the challenge of occlusion and sparsity.
-To demonstrate this, we perform the relax the criterion of eligibility of ground truth such that they only need to contain at least one point emitted by any agents in the V2X network.
-The performance shown in Tab.VI indicates that on a larger set of ground truths, Late Collaboration does improve performance, compared to single-vehicle (no collaboration).
+Remarkably, our collaboration approach using MoDAR points outperforms the early collaboration by $4.7\%$ (3.35 mAP).
+This reaffirms our design philosophy that a good multi-agent collaborative perception framework can be made on the foundation of a good single-agent perception model and a simple collaboration method.
+Enhance Detection of Completely Obstructed Regions Objects
+Besides enabling the ego vehicle more accurately detect objects that are visible to itself, another great benefit of collaborative perception is to help the ego vehicle see the objects that are completely obstructed - those that do not contain any points of its point cloud $\mathcal{P}_{e}^{t}$, thus overcoming the challenge of occlusion and sparsity.
+To demonstrate this, we relax the criterion of eligibility of ground truth such that they only need to contain at least one point emitted by any agents in the V2X network.
+In this experiment, we add a strong mid-collaboration method - DiscoNet [3] to the set of baselines to showcase the capability of our approach compared to the state-of-the-art.
+In the framework of DiscoNet, connected agents, which have identical detection architecture, exchange BEV images of their point cloud.
+The ego vehicle fuses the BEV image of its point cloud with the exchanged BEV via a fully connected collaboration graph to use as the input to the detection head.
+In addition to the detection precision, we also measure the bandwidth consumption of each collaboration method.
+Our measurement is based on the raw uncompressed form of the exchanged data to facilitate a fair comparison.
+The results shown in Tab.VI indicate that on a larger set of ground truths, Late Collaboration does improve performance, compared to single-vehicle (no collaboration).
 The improvement is significant (at least $8.35$ mAP) even in the poorest setting where the set of exchanged detection $\left\{\mathcal{B}_{i}|i\in\{0,2,3,4,5\}\right\}$ is $0.2$ seconds behind the time of query, thus resulting a spatial misalignment between detected objects that are exchanged and their associated ground truths if the underlying objects are dynamic.
-This can be explained by a significant number of static and slow-moving objects present in an intersection whose past detections remain true positive at the present.
-When detections of dynamic objects are better accounted for as in the sync setting where agents exchange their detections at the same timestamp as the query, the performance is largely improved by $9.29$ mAP, compared to the async setting.
+This can be explained by a significant number of static and slow-moving objects present in intersections whose past detections remain true positives at the present.
+When detections of dynamic objects are perfectly accounted for as in the sync setting where agents exchange their detections at the same timestamp as the query, the performance is largely improved by $9.29$ mAP ($15.2\%$), compared to the async setting.
 However, this sync setting is unrealistic because different agents have different detection rates.
 Interestingly, propagating detections using scene flow as in the async prop.
-setting can reach $96.20\%$ the performance of the sync setting ($2.68$ mAP behind).
+setting can reach $96.2\%$ the performance of the sync setting ($2.68$ mAP behind).
 This implies the effectiveness of scene flow estimated by our Aligner++.
-In this setting, our collaboration method using MoDAR no longer outperfroms the early collaboration baseline as some distant objects (with respect to the ego vehicle) are inevitably missed by connected agents, thus remaining invisible to the ego vehicles and being unable to be detected.
-However, our method still maintain favorable performance of $98.2\%$ of the early collaboration’s while attaining as low bandwidth usage as late collaboration.
-TABLE VI: performance of collaboration methods on ground truth that are visible to at least one agents
-Collaboration Method mAP
-None 52.84
-Late - async 61.19
-Late - async prop. 67.80
-Late - sync 70.48
-Early 78.10
-Ours - MoDAR 76.72
-Ours - MoDAR 75.41
-and Foreground
+TABLE VI: performance of collaboration methods, measured by mAP, on ground truth that are visible to at least one agent
+Collaboration Sync Async Bandwidth
+Method Usage (MB)
+None 52.84 - 0
+Late 70.48 61.19 0.01
+67.80  [^1] Exchanged detections are propagated using scene flow according to (5) before fusion by Non-Max Suppression.
+Mid - DiscoNet  [^2] Our implementation based on [3] 78.70 73.10 25.16
+Early 78.10 77.30 33.95
+Ours 79.20 76.72 0.01
+In the sync setting, DiscoNet narrowly exceeds Early Collaboration by $0.6$ mAP ($0.7\%$), thus showing the benefit of the deep interaction among agents powered by their dense collaboration graph.
+However, this benefit comes at the cost of relatively high bandwidth usage of $25$ MB per broadcast BEV image on average which is only $25\%$ less than the average bandwidth required by Early Collaboration.
+The operation of our method in this setting is as follows: MoDAR points broadcast by other agents have the same timestamp as the point cloud of the ego vehicle; therefore their propagation using scene flow is skipped.
+Our method slightly exceeds DiscoNet by $0.5$ mAP ($0.6\%$) to reach the highest performance while consuming the least bandwidth among collaborative methods.
+This confirms the effectiveness of MoDAR in facilitating collaboration among connected agents.
+In the async setting, DiscoNet is severely affected by the misalignment among exchanged BEV images caused by timestamp mismatch as its performance dropped by $5.6$ mAP ($7.1\%$).
+Our method is affected because the misalignment between exchanged MoDAR points and their corresponding ground truth is explicitly taken into account by the propagation of MoDAR points using scene flow.
+As a result, the performance gap between ours and DiscoNet is extended to $3.62$ mAP ($5\%$).
+Early Collaboration shows the most robustness against timestamp mismatch with an $0.8$ mAP ($1\%$) drop.
+This can be explained by the detector’s ability to implicitly handle to some extent the misalignment among point clouds.
+This phenomenon is well known in the literature on multi-frame single-vehicle object detection [6, 38, 8, 39] where models trained on the concatenation of 0.5-second point cloud sequences enjoy a $30\%$ mAP gain compared to models trained on individual point clouds.
+Support Heterogeneous Detector Networks
+To show that our collaboration method supports heterogeneous networks out of the box, we perform an experiment where we start with a network of agents uniformly using PointPillar as their detectors and gradually replace PointPillar with SECOND in one agent after another.
+The result of this experiment is shown in Fig.10.
+Figure 10: Performance of the ego vehicle in heterogeneous collaboration networks made PointPillar and SECOND detectors.
+As the number of SECONDs increases from 0 to 6 - the size of the network, the ego vehicle’s performance transits from its value in a full-PointPillar network to its value in a full-SECOND network without catastrophe dropping out of the lower bound of Late Collaboration.
+As the number of SECOND increases, the performance of the collaborative perception gradually transits in a downward trend from $76.72$ mAP of a full-PointPillar network to $71.94$ mAP of a full-SECOND network, without dropping below the lower bound of Late Collaboration, which is the case of mid-collaboration method without dedicated modules for bridging the domain gap [5].
+This downward trend is because an individual PointPillar achieves higher AP than an individual SECOND on the V2X-Sim dataset when being trained under the same setting.
+More interestingly, the combination made of the RSU using SECOND and all connected vehicles using PointPillar archives the best performance of $77.74$ mAP, reaching $99.5\%$ performance of Early Collaboration in sync setting.
+Relation between Performance and Network Size
+Finally, to study how the performance of collaboration perception evolves with respect to the number of agents in the V2X networks, we gradually increase the number of participants in the collaboration.
+As can be seen in Fig.11, the collaboration with only the IRSU brings $15.09$ mAP ($28.6\%$) improvement.
+The magnitude of performance gain reduces and eventually saturated as the number of agents increases.
+An increased number of agents results in higher overlap in agents’ field of view which makes objects in the overlapped regions well detected by many agents.
+However, this overlap also means the field of view of the collaboration network as a whole does not increase proportionally with the number of agents.
+Moreover, exchanging detected objects made by distant agents may not get used by the ego vehicle because they are outside the predefined detection range.
+Qualitative
+The qualitative performance in Fig.12 shows that collaborative perception using MoDAR points enables the ego vehicle to detect objects that are occluded or have a few to zero LiDAR points due to long range.
+Particularly, the vehicle at the top of Fig.12a is occluded with respect to the ego vehicle due to the presence of a large vehicle.
+However, this occluded vehicle is successfully detected thanks to a single MoDAR point produced by the IRSU.
+Another example of occluded vehicles that are successfully detected is Fig.12e where vehicles in the bottom right corner are occluded by a building-like structure.
+The illustration of successful detection at long range thanks to MoDAR points can be found in Fig.12b, Fig.12d, and Fig.12f where vehicles near their edges have a few or even zero LiDAR points of the ego vehicle.
+Figure 11: The performance of ego vehicle’s detection increases sharply in collaboration with only IRSU, then slows down and saturates at $98.2\%$ of Early Collaboration in sync setting as the number of connected agents increases.
+(a) 1 RSU & 3 CAVs
+(b) 1 RSU & 1 CAVs
+(c) 1 RSU & 3 CAVs
+(d) 1 RSU & 2 CAVs
+(e) 1 RSU & 3 CAVs
+(f) 1 RSU & 1 CAVs
+Figure 12: Qualitative performance of our method on three samples of the V2X-Sim dataset.
+The number and type of connected agents are annotated in the title of each image.
+Blue points are LiDAR points collected by the ego vehicle.
+Gray points are LiDAR points collected by other agents which are displayed for visualization purposes only.
+Orange stars denote the MoDAR points broadcast by other connected agents.
+Green solid and dashed rectangles respectively represent ground truth visible and invisible to the ego vehicle.
+Red rectangles are the detections made by the ego vehicle using our method.
+V Conclusions and Future Work
+In this paper, we develop a collaborative perception framework based on the foundation made of good single-agent perception models and an effective late-early collaboration method.
+The simplicity of our collaboration method facilitates its practicality by (i) minimizing bandwidth usage, (ii) dismissing the need for inter-agent synchronization, (iii) making minimal changes to single-agent object detectors, and (iv) supporting networks of heterogeneous detectors.
+Importantly, this practicality does not compromise the overall performance as our collaboration method exceeds Early Collaboration on the V2X-Sim dataset.
+The success here paves the way for our next development which is the demonstration of our method on a real-world V2X network.
+VI Acknowledgement
+This work was granted access to the HPC resources of IDRIS under the allocation 2021-AD011012128R1 made by GENCI.
+This work has been supported in part by the ANR AIby4 under the number ANR-20-THIA-0011.
+This work was carried out in the framework of the NExT Senior Talent Chair DeepCoSLAM, which is funded by the French Government, through the program Investments for the Future managed by the National Agency for Research (ANR-16-IDEX-0007), and with the support of Région Pays de la Loire and Nantes Métropole.